### v1

# Common Corpus: The Largest Collection of Ethical Data for LLM Pre-Training

Abstract
Large Language Models (LLMs) are pre-trained on large amounts of data from different sources and domains. These data most often contain trillions of tokens with large portions of copyrighted or proprietary content, which hinders the usage of such models under AI legislation. This raises the need for truly open pre-training data that is compliant with the data security regulations. In this paper, we introduce Common Corpus [^1] https://huggingface.co/datasets/PleIAs/common_corpus, the largest open dataset for language model pre-training. The data assembled in Common Corpus are either uncopyrighted or under permissible licenses and amount to about two trillion tokens. The dataset contains a wide variety of languages, ranging from the main European languages to low-resource ones rarely present in pre-training datasets; in addition, it includes a large portion of code data. The diversity of data sources in terms of covered domains and time periods opens up the paths for both research and entrepreneurial needs in diverse areas of knowledge. In this technical report, we present the detailed provenance of data assembling and the details of dataset filtering and curation. Being already used by such industry leaders as Anthropic and multiple LLM training projects, we believe that Common Corpus will become a critical infrastructure for open science research in LLMs.
1 Introduction
Large Language Models have been defined by large training data. While there are several candidates for the first modern language model based on transformer architecture, including GPT-1 (Radford et al., ), ULMFIT (Howard and Ruder, 2018), or Sentence Neuron (Radford et al., 2017), it is commonly acknowledged that “large” models start with GPT-3 (Brown et al., 2020). Requiring a corpus of 300 billion tokens, GPT-3 introduced a standard training data pipeline shared by nearly all language models to date: large-scale processing of web datasets (45 TB of compressed source data from Common Crawl) and additional digitized sources (e.g. Books3). Until 2025, LLM training data has grown on a logarithmic curve. The latest generation of publicly documented language models including DeepSeek v3 (DeepSeek-AI, 2025), Gemma 3 (Google, 2025), Llama 4 (Meta AI, 2025) or Qwen 3 (Qwen, 2025) have been trained on 14-36 trillion tokens. Even the gaining popularity sub-category of small language models (Wang et al., 2025) relies on large amounts of training data to fit scaling laws: Qwen 3 0.6B was trained on 36 trillion tokens, about 3,000 times more data than recommended by the original Chinchilla laws (Hoffmann et al., 2022).
As data curation also became a primary concern, the collection, maintenance, processing, and filtering of data became one of the primary cost of standard language model training. This does not factor in even larger hidden costs: the negative externalities affecting competing markets, the digital commons, and society at large.
While data scraped from the web is publicly available, it is not always in the public domain. Most web data do not have sufficient metadata to determine whether it is permissively licensed. NLP practitioners have relied on the protection of fair use, claiming that the transformative nature of the use of the data allows them to use this data to train language models. There are increasingly more legal challenges to the use of this data. The New York Times sued Open AI for copyright infringement, alleging that Open AI trained their models on their articles (Roth, 2023, Pope, 2024). Due to concerns about indirect commercial exploitation, many rightholders have implemented either hard technical measures or legal provisions against model training. In 2024, it was estimated that for Terms of Service crawling restrictions, a full 45% of C4 is now restricted (Longpre et al., 2024b) and 5% is fully blocked for scraping with a disproportionate impact over quality sources (Longpre et al., 2024b). Restrictions not only affect language model pre-training but also the quality of search engine indexation and a variety of research projects analyzing and collecting content at scale. Even projects dedicated to knowledge access have faced significant pressure from AI crawlers and implemented protections that negatively impact access and user experience.
These legal uncertainties have significantly impeded the development of open science research on LLMs. Previously reproducible research artifacts have been removed or taken down, impacting pre-training data, continuous pre-trained models, and evaluation datasets. Books3, which has been used in datasets like the Pile (Gao et al., 2020) and by multiple AI labs, including Meta and Microsoft, faced legal challenges (Brittain, 2023), and the original dataset was ultimately removed (Van der Sar, 2023). The LAION dataset was demonstrated to contain CSAM (Birhane et al., 2021, Thiel, 2023), and then was taken down (LAION, 2023). Later, it was re-released once suspected CSAM was removed (LAION, 2024). The Dutch-focused model GEITje was taken down (Rijgersberg, 2025) due to complaints about it having been trained on the Dutch Gigacorpus, in order to avoid legal disputes. Finally, the widely used benchmark, the Mathematics Aptitude Test of Heuristics (MATH) dataset (Hendrycks et al., 2021), was removed from Hugging Face via a DMCA takedown. All of these artifacts, which were released to further open development and evaluation of language models, were removed suddenly, making previous work unreplicable. These takedowns and legal challenges also represent a sizeable loss of investment for developers, who are often independent or small research organizations.
In part, as a reaction to the use of publicly available but not permissively licensed data, web text is becoming increasingly more difficult to acquire and use. In an analysis of popular datasets such as C4 (Raffel et al., 2020), RefinedWeb (Penedo et al., 2023), and Dolma (Soldaini et al., 2024), Longpre et al. (2024c) found that just in the last year, 5% of all tokens in C4 now have restricted use, with a disproportionate number of those tokens coming from the best-maintained, most critical sources. This is largely due to changes in content owners’ and hosts’ preferences, which are changing to no longer allow scraping, especially for the purposes of training AI models.
Since 2024, several initiatives have emerged to collect open data with clear licensing and in the English language. This includes: C4C, Open License Corpus, a 228 billion token corpus from a mix of public domain texts and open source code under free licenses not requiring attribution (Min et al., 2024), KL3M, a 1.2 trillion tokens corpus of administrative texts and structured data mostly from the US federal public domain (Bommarito et al., 2025), Common Pile, an ongoing data collection of 1 trillion tokens from a variety of recent sources, including a filtered version of Common Crawl data (Creative Commons Common Crawl) (Kandpal and Raffel, 2025). These projects are all monolingual, restricting in effect the reach of open-data language models to the English-speaking audience. In contrast, the most ambitious multilingual collection of permissive content pre-dates Large Language Models: C4C (2016), containing 12 million web pages in more than 50 languages filtered by Creative Commons Licenses (Habernal et al., 2016).
In this paper, we present Common Corpus, which has been an ongoing effort since 2024 (see Appendix A for the development history). It has grown to become the largest open dataset for pre-training Large Language Models at about 2 trillion tokens and the only one in its size range available in multiple languages. Thus, Common Corpus is a significant contribution to the AI Commons. Through the release of this dataset, we show that open LLM research and development is possible while meeting legal and regulatory requirements—even in the EU, where AI regulations are currently the strictest. Here, we detail the composition of Common Corpus (Section 3), as well as the data provenance, license verification, and further processing (Section 4). Next, we describe the early impact of Common Corpus on the AI community (Section 5) and highlight the limitations of the work that has been done at the time of publication. Despite its size, Common Corpus is still far from covering the entire range of available resources for pre-training commons: we attribute this discrepancy to an open data paradox as major sources of open content are paradoxically little visible online and even more so in the leading sources currently used for pre-training. By describing the unique challenges coming with the aggregation of large open source, we aim to inspire further responsible data collection initiatives.
Figure 1: A schematic world map of languages in Common Corpus with a log-scaled distribution of document counts. For each language, we chose a city that is located in the region where this language is most specific to. To avoid outliers, we show only languages with 1000+ documents.
2 About Common Corpus
When talking about Common Corpus data, we use the word “open” in the strongest possible sense. Not only is the data available, but we also provide essential details about the data provenance, data processing, and important information about the contents of each dataset. Following the Open Source Initiative’s definition of open-source AI, our data is open in terms of openness of use, meaning that use is permitted for “any purpose and without having to ask for permission” (Open Source Initiative, 2024). To achieve this, models must be trained on datasets that are free from copyright or other relevant legal limitations. This is currently a limitation of existing open datasets for training LLMs.
Common Corpus, therefore, provides valuable training tokens that will not be subject to the same restrictions. Additionally, the data in Common Corpus are different from other corpora, primarily composed of web text. Common Corpus contains multilingual data in a variety of high- and low-resource languages (see Figure 1 for language distribution), covering diverse genres, time periods, and domains (in Section 3, we detail each part of the dataset). Therefore, Common Corpus contributes to data diversity in the open pre-training data ecosystem. This is important for developing powerful and generalizable model performance. Common Corpus can be used on its own or in conjunction with existing open datasets, according to one’s needs and the desired use case of a language model.
(a) A timeline of the main collections with their numbers of documents in the Common Corpus.
(b) A two-component t-SNE visualization of a subset of the Common Corpus.
Figure 2: Temporal and semantic overview of the Common Corpus collections.
Dataset Documents Words Tokens
Open Government 074,727,536 0,257,233,670,261 0,406,581,454,455
Open Culture 093,156,602 0,549,608,763,966 0,885,982,490,090
Open Science 019,220,942 0,147,305,783,453 0,281,193,563,789
Open Code 202,765,051 0,077,669,169,092 0,283,227,402,898
Open Web 096,165,348 0,033,208,509,065 0,073,217,485,489
Open Semantic 030,072,707 0,023,284,201,782 0,067,958,671,827
Other 000,925,462 0,000,328,160,421 0,000,486,099,734
Total 517,033,648 1,088,638,258,040 1,998,647,168,282
Table 1: Dataset composition of Common Corpus. For each collection, we report the total number of documents, words (whitespace-separated), and tokens.
Common Corpus was developed with consideration for ongoing conversations about best practices for open-source LLM development (The AI Alliance, 2024, Longpre et al., 2024a, Duprieu and Berkouk, 2024, Baack et al., 2025). As PleIAs was one of the contributors at the Dataset Convening, we highlight our adherence to the best practices that were suggested by Baack et al. (2025):
  • Provide useful documentation. We provide information about dataset provenance and processing (Sections 3 and 4) and share key statistics to help potential users understand the applications of the dataset. This is crucial, as dataset documentation improves reproducibility, helps prevent misuse, and aids downstream users to best utilize the dataset (Longpre et al., 2024a).
  • Follow and record preference signals. In the metadata, we include the source URL and license information for the vast majority of the corpus.
  • Increase diversity and involve local communities to identify relevant data sources. This dataset includes data from a variety of languages, coming from high-quality sources, and the multilingual part was never machine-translated.
  • Share advancements to foster reciprocity and give back. In addition to the dataset, we release many of the tools we developed in order to create the final dataset (Section 4).
  • Do not use openly licensed data without regard for its quality or fitness for purpose. In particular, for the dataset in the public domain, we engage in extensive OCR correction and toxicity filtering in order to bring datasets up to standard (Section 4).
  • Do not capture highly sensitive data. We remove personally identifiable information from our datasets (Section 4.4).
Common Corpus aims to support the pre-training of fully open and auditable LLMs by making it legal to release the source even without the provision of fair use. It has been used to create a wider range of language model artifacts, including multimodal datasets, classifiers, synthetic datasets, and benchmarks. Beyond the main dataset, Common Corpus works as an open science infrastructure dedicated to the entire lifecycle of language models. As defined by UNESCO, it is a shared research infrastructure that is needed to support open science and serve the needs of different communities (Unesco, 2021). We argue this is the first point in time where there has been sufficient knowledge and infrastructure to collect and clean a dataset on this scale, which meets the legal and ethical criteria we have outlined.
2.1 Composition
Common Corpus is available on HuggingFace as an aggregation of 10,000 parquet files and is composed of six collections: Open Government, Open Culture, Open Science, Open Web, Open Code, and Open Semantic. In total, the number of tokens in Common Corpus is 1,998,647,168,282. The token counts [^2] We report token counts in terms of the Pleias base tokenizer https://huggingface.co/PleIAs/Pleias-350m-Preview trained on a subsample of Common Corpus. in each collection are listed in Table 1. We visualize the timeline of the collected documents and embeddings of a subsample in Figure 2. Each collection is composed of multiple datasets, for which we provide details about provenance and other key information in the corresponding subsections. Each data object contains a license, language(s), a collection/domain of specialization, and other metadata, allowing one to filter out a desired subset.
Language Tokens
English 968,757,721,747
French 275,358,437,630
German 112,127,458,251
Spanish 046,514,142,421
Latin 036,031,591,540
Italian 024,681,637,575
Polish 012,146,688,669
Greek 011,376,498,056
Portuguese 010,262,747,943
Russian 009,439,453,633
Table 2: Token counts for the ten most represented languages in Common Corpus.
License type Tokens
Public Domain 1,138,508,375,958
$C$-By 0,287,749,264,457
MIT 0,142,694,227,607
$C$-By-SA 0,074,768,060,836
Apache-2.0 0,068,750,977,037
BSD-3-Clause 0,018,483,944,333
Open license 0,010,432,513,767
BSD-2-Clause 0,005,497,145,480
$C$-BY-4.0 0,002,110,966,243
CC0-1.0 0,001,877,206,195
Table 3: Token counts for the ten most common licenses in Common Corpus.
The majority of the data in Common Corpus is in the public domain (see Table 3). The license for each document is provided in the metadata, so the dataset can be easily filtered by license as needed for a particular use case.
Common Corpus is multilingual (see Figure 1 and Table 3 for language coverage). Unlike many pretraining mixes, which are coposed mostly of English and code data, our dataset has strong representation for several languages, with the top nine languages constitute at least 10B tokens each [^3] Language distribution was computed using the fastText language identification model.. While we highlighted some of the issues faced in making open datasets above, all of these issues are exacerbated for languages other than English. Even in relatively high-resource languages like French, these problems are compounded by the fact that there is much less data available, and most tools generalize poorly to languages other than English. Additionally, Kreutzer et al. (2022) showed that many multilingual datasets contain a lot of low-quality or entirely unusable data. Many of the datasets they analyzed contained less than 50% of usable text, with 15 sources containing no usable data at all.
3 Provenance
In this section, we present the details about collections that comprise the Common Corpus, accompanied by the detailed information about the data sources, token count breakdowns, and the main included languages in Appendix B.
3.1 Open Government
Open Government is a set of financial, legal, and administrative data in the public domain. In total, the dataset contains more than 406B tokens and comprises two main datasets: Finance Commons and Legal Commons. See Appendix B.1 for detailed data composition.
Finance Commons.   This is the largest collection of financial documents in the public domain, comprising more than 14 billion words (more than 23 billion tokens). The documents come from a wide time range, from the mid-20th century all the way to 2024. Like many of our other datasets, Finance Commons is also multilingual. Most of the documents are in English, French, and German, but there are also texts in languages such as Romanian, Bulgarian, and Latvian. Additionally, this is a multimodal dataset. It includes more than 1.36 million original PDF documents from AMF and the WTO. The documents constitute a wide coverage of in-house layouts and formats produced by industrial and economic sectors. This makes this dataset ideal for developing the next generation of open-data multimodal models. One application for this dataset is to develop vision-language models (VLMs) for advanced document segmentation and processing. These documents also contain vast amounts of structured data, which is also a promising area of research that Finance Commons can help drive forward.
Legal Commons.   This is a collection of legal and administrative datasets. The datasets come mostly from the EU and the US and cover a wide range of languages. These datasets are useful for developing language models with legal knowledge, as well as models that are ideal for document processing in official administrative applications.
3.2 Open Culture
Open Culture is an aggregation of vast cultural heritage datasets containing both monographs and periodicals for over 13 languages: French, English, German, Spanish, Portuguese, Italian, Dutch, Luxembourgish, Danish, Swedish, Serbian, Czech, and Greek. There are also small portions of data in a wide variety of other languages, including Arabic, Bengali, Latin, Persian, Russian, Sanskrit, and Urdu.
A large part of Open Culture is compiled from Collections As Data (CAD) — large dumps of texts, datasets, PDFs, and even raw XML output (METS/ALTO). CAD initiatives, thus, considerably simplify dataset aggregation and are a major contribution to the digital commons ecosystem. All other parts of Open Culture have been collected on a resource-by-resource basis using APIs and other standard retrieval methods whenever available. The largest extractions of this kind include Internet Archive (about 2 million monographs in multiple languages) and Delpher (50,000 Dutch monographs and periodicals filtered to match the Dutch copyright law for public domain). We managed to compile a large multilingual collection despite such challenges, as poor OCR quality, which we partly solved through the development of OCR correction tools (see Section 4), text segmentation issues, and sometimes irrecoverable deterioration of the original support. For the detailed dataset composition, refer to Appendix B.2.
All Open Culture documents are in the public domain, which means their copyright has expired after a given term and there are no limitations on their reuse. For certain content, or in cases where we could not rely on the guarantee of established cultural heritage institutions, we implemented our own internal rights verification process. This process follows specific criteria, including author life and data object creation time, and takes into account that we only collected cultural heritage content from institutions based in the US or the EU (see the complete criteria list in Appendix C).
Open Culture data is also rich from a cultural and stylistic standpoint and can be used to train multilingual language models with more diverse and creative writing styles. As LLMs are trained on extremely large corpora to maximize next-word prediction accuracy, LLM-generated text can often lack in personality and be boring or generic (Jones and Bergen, 2024). This feature of language models stands in contrast with one of their most common uses. In an analysis of WildChat (Zhao et al., 2024), a dataset of 1 million user interactions with ChatGPT, Longpre et al. (2024c) found that over 30% of user requests involved creative compositions such as fictional stories, role-play, or poetry generation. At the same time, creative writing is poorly represented among datasets used to train LLMs, which mainly comprise web text (Longpre et al., 2024c). Therefore, Open Culture contributes data that can be used to train models for creative writing without violating copyright law. In addition, as many of the Open Culture datasets are historical (coming from the 18th-19th centuries, or even earlier; see Figure 2(a)), this collection also enables the development of historical language models. The metadata includes document creation year, which enables researchers to develop language models with a cutoff of the training data creation date.
3.3 Open Science
The Open Science collection includes scientific papers and other documents (theses, book reviews, clinical trials, etc). Following on the development of a global open access movement, these documents have been made increasingly available in open archives (preprints) or directly through open science publishers and infrastructure. Scientific content has become a primary focus for training data curation, due to its impact on performance on reasoning and advanced world knowledge tasks. Yet, the lack of licensing information has until now partly hindered reuse. The Semantic Scholar Open Research Corpus from Allen AI includes 81.1 million articles in English under an Open Data Commons Attribution License, allowing for the free reuse of the aggregated metadata while still acknowledging the remaining copyright of individual authors (Lo et al., 2020). The Pile incorporated dumps made available by arXiv and PubMed Central, also exclusively in English (Biderman et al., 2022). Finally, the BigScience project assembled several curated multilingual scientific datasets like the French HAL as part of the training data for BLOOM (Workshop et al., 2023).
The Open Science collection was made possible largely due to the recent development of OpenAlex [^4] https://openalex.org/, the largest open catalogue of scientific documents. OpenAlex maintains an expansive API search engine tracking detailed metadata for each indexed item, including the licensing, as well as a link to the original resource, which is generally in PDF format. We filtered OpenAlex on the three following licenses: CC-By, Public Domain/CC0, and CC-By-SA. The largest share of resources is available under CC-By, which is currently the recommended license by the Open Access definition. Open Science also includes smaller subsets, such as a direct extraction of arXiv articles available in CC-By and some European-specific resources not currently well indexed on OpenAlex (the exact distribution of token counts can be found in Appendix B.3).
Due to the specificity of open scientific publishing, the Open Science collection has less linguistic diversity, with nearly 85% of documents currently available in English.
3.4 Open Code
The Open Code collection comprises code data under a vast variety of free licenses, which allows NLP practitioners to train models on public domain code for either coding applications or in order to improve certain model performance on natural language reasoning, world knowledge tasks, mathematics, and structured output tasks (Aryabumi et al., 2024, Petty et al., 2024, MA et al., 2024). The code data we use comes from the Stack v1 and v2 (Kocetkov et al., 2023, Lozhkov et al., 2024). The Stack v1 contains 6.4TB of data and covers 30 programming languages, while the Stack v2 is approximately ten times bigger at 67.5TB and covers over 600 programming languages. All the code data is made available with a direct link to the original resource on GitHub. In total, Open Code contains 283,227,402,898 tokens (see most common languages in Appendix B.4).
[…]
3.5 Open Web
In accordance with the general focus of Common Corpus on curated content, the Open Web collection currently includes four major web sources:
Wikipedia and Wikisource.   Wikimedia projects have always been major sources for language model training due to their reliability, extensive coverage, and textbook-like style. Despite this centrality, there is still a range of unresolved challenges with the most common versions available for training. The raw source of Wikimedia projects is made available in a specific mediawiki syntax, including a lot of project-specific models, tags, and conventions. The parsing of models is especially not straightforward, as they can either format existing text or remove or include external content (transclusion). As part of Wikimedia Enterprise, the Wikimedia Foundation created entirely new dumps from the rendered HTML sources, which in effect ensure that they include all the text made available to readers.
Youtube Commons.   YouTube Commons is our collection of audio transcripts of 2,063,066 videos uploaded on YouTube under a standardized CC-By license [^5] https://huggingface.co/datasets/PleIAs/YouTube-Commons.
StackExchange.   This is a collection of user-generated forums and Q&A made available under the CC-By-SA license. We reused the version from The Pile (Biderman et al., 2022).
A major objective will be the integration of web archives filtered by permissive license. Since 2016, several projects have attempted to reidentify Creative Commons licenses from web archives at scale including C4C (multilingual) (Habernal et al., 2016) and more recently CCCC (from Allen AI - in English) and most recently Common Crawl Creative Commons Corpus (C5; the first multilingual C4 derivative) [^6] https://huggingface.co/datasets/BramVanroy/CommonCrawl-CreativeCommons All theses projects struggled with license identification. While license mentions are frequently normalized with a direct link or logo to Creative Commons, there is no guarantee they really concern the entire content: “a blog page contains many photos and each photo is licensed under different CC-license type or a blog home page with many articles and each article is licensed under different CC-license type.”(Habernal et al., 2016). We hope this limitation could be overcome by a combination of web domain curation and fine-grained curation and annotation by a language model.
3.6 Open Semantic
Semantic data is the latest set added to Common Corpus and currently includes only one collection: Wikidata. First created in 2011, Wikidata hosts 100 million documented items and several billion factual statements encoded as RDF triples. It has grown to become a critical web infrastructure, used by Google for search disambiguation and currently embodying Tim Berners-Lee’s ambitious vision for “a web of data”. Despite the rising interest in mixed LLM/knowledge graph methods, Wikidata has hardly been used in language models. The largest initiative to date is Kelm, a collection of 15 million synthetic sentences generated by Google from English-speaking statements (Agarwal et al., 2021). A persistent challenge has been the exclusive availability of Wikidata dumps under formats optimized for data exchange rather than language model training.
Thanks to a collaboration with Wikimedia Deutschland, the entire set of Wikidata has been adapted in natural language and added to Common Corpus. This is to date the only available textual collection of Wikidata covering the entire range of 300 languages. Data processing involved the translation of items and properties into formal language sequences as simple natural language sequences, without textual synthesis: “Q41309 | P:27 | Q171150” becoming “Franz Liszt country of citizenship Kingdom of Hungary”. Within each entry, we provide all the available translations as consecutive blocks separated by a newline, anticipating that this may contribute to language alignment.
4 Cleaning and Curation
4.1 Text Segmentation
For text segmentation, we developed Segmentext [^7] https://huggingface.co/PleIAs/Segmentext (see example in Appendix D.1). It was trained to be resilient to broken and unstructured texts with digitization artifacts and ill-recognized layout formats. Given the diversity of the training data, Segmentext should work correctly on diverse document formats in the main European languages.
4.2 OCR Error Detection
We developed two pipelines to determine the digitization quality of different datasets in order to determine the appropriate treatment before datasets can be used for pre-training:
OCRoscope [^8] https://github.com/Pleias/OCRoscope.    This is a tool based on the language identification tool cld2999https://pypi.org/project/pycld2/. As cld2 cannot attribute non-recognized strings to any language, OCRoscope works by measuring the number of non-recognized 7-grams. OCRoscope provides a standardized rate of OCR quality and rate of non-character content for at least 80 languages. An example of OCPoscope work is presented in Appendix D.2. Complete processing by OCRoscope yielded a rate of 41% non-recognized 7-grams, which results in an OCR quality rate of 59%. In comparison, the self-estimated rate of OCR valid words by the French National Library is significantly higher (85%) for the whole document.
OCRerrcr [^10] https://huggingface.co/PleIAs/OCRerrcr.   This is a DeBERTa-v2-style language model with 400M parameters. OCRerrcr achieves higher accuracy than OCRoscope, but is more computationally intensive.
4.3 OCR Correction
We developed OCRonos [^11] https://huggingface.co/PleIAs/OCRonos model based on Llama 3 8B (Grattafiori et al., 2024). OCRonos is versatile and supports the correction of OCR errors, wrong word cuts or merges, and overall broken text structures. The training data includes a highly diverse set of OCR-ed texts in multiple languages, mostly coming from uncorrected versions of Open Culture and Open Government. On highly deteriorated content, OCRonos can act as a synthetic rewriting tool rather than a strict correction tool. An example of OCRonos work is presented in Appendix D.3. OCRonos contributes to making challenging resources usable for LLM applications and, more broadly, search retrieval. It is especially fitting in situations where the original PDF sources are too damaged for correct OCR or are not possible to retrieve.
OCRonos is generally faithful to the original material, provides sensible restitution of deteriorated text, and will rarely rewrite correct words. A common issue with OCR correction has been language switching: due to the inherent noise in the input text, an LLM will transcribe in a different language or script. The issue has been especially observed in smaller generalist models like GPT-3.5 or Claude-Haiku. OCRonos largely mitigates this issue.
4.4 PII Removal
Personally Identifiable Information (PII), which is any information that can be used to distinguish or trace an individual’s identity, is protected under legislation such as GDPR. Consequently, the new reguations put restrictions on LLM training data. In an analysis of large open datasets, there is a staggering amount of personal data in widely used datasets, e.g., large quantities of phone numbers in RedPajama, email addresses in S2ORC and peS2o, and IP addresses in the Stack (Elazar et al., 2024).
To identify and replace PII, we use Microsoft’s Presidio [^12] https://microsoft.github.io/presidio/, an open-source state-of-the-art tool. With the base settings, Presidio identified on average 55-60% of texts that included phone numbers due to different possible number formats. By applying custom regular expression patterns that include most phone numbers, we increased this accuracy to 85%. Typical methods of handling PII include removing it, replacing it with tags, and partial anonymization. These transformations substantially alter the format of PII, which could undermine the model’s understanding of the text or interfere with its ability to process text with real PII. Instead, we replace PII with fictuous but realistic values.
4.5 Toxicity Detection
In addition to posing legal and regulatory issues, web data is a major source of harmful and biased content (Common Crawl was shown to contain sexual content, hate speech, and racial and gender biases (Luccioni and Viviano, 2021)) and often suffers from low-quality and machine-generated text (Dodge et al., 2021). Public Domain data, e.g., in Open Culture, do not pose the same legal challenges but introduce new ones. Many texts there are historical periodicals and monographs from at least 80 years ago. Cultural norms surrounding the discussion of certain ethnic groups, women, and themes such as violence have changed dramatically. Many of these texts, therefore, do not meet modern ethical standards. Training models on these texts would lead to the reproduction and circulation of harmful language.
[…]
5 Impact
As early as August 2024, the first version of Common Corpus was highlighted as a “massive” part of pre-training data commons in European languages, especially contributing to making French a high-resource language (Ali and Pyysalo, 2024). In February 2025, Common Corpus was acknowledged as one of the main deliverables of the Paris AI Summit, as the “largest open database to train Large Language Models” [^15] https://www.elysee.fr/admin/upload/default/0001/17/373d6bcd8ea7f84a701c050bb9cab ccf3ed95c2b.pdf.
Common Corpus has already been incorporated in a wide range of LLM projects. As of May 2025, seven European SLMs and LLMs have been trained on at least some part of Common Corpus, and more will come throughout the year 2025. Among these are the models in the Pleias 1.0 family, including three SLMs (350 million, 1.2 billion, and 3 billion parameters), which were exclusively trained on Common Corpus. As such, they are fully reproducible in Europe. They were trained on the Jean Zay supercomputer in France (using H100 GPUs, under a Grand Challenge compute grant), and, specifically for the 1.2B model on the TractoAI’s cloud platform [^16] https://tracto.ai/. They come with a new tokenizer with 65,536 unique tokens, also trained on a representative sample of the Common Corpus for increased training efficiency and better multilingual support. During training, we experimented with different data mixtures of Common Corpus and derived a “refined” version for late training of about 1 trillion tokens, with more intensive, curation-based filtering. Two models derived from Pleias 1.0, Pleias-RAG-350M and Pleias-RAG-1B, are currently leading evaluations in their size range for retrieval augmented evaluation. Part of the performance can be attributed to the solid coverage of the main European language in Common Corpus, as well as a very extensive proportion of PDF-extracted sources that are very close to content commonly processed in the RAG production setting.
In January 2025, Barcelona Supercomputing Center released the Salamandra series of models under three different sizes: 2, 7, and 40 billion parameters (Gonzalez-Agirre et al., 2025). A major objective was enhanced support for a wide range of European languages, including lower-resource languages. The French part of Open Culture, French-PD, was incorporated into the overall data mixture. It was the main French dataset and the second-largest curated source. A French 7 billion parameters SLM, Lucie, incorporated a wider range of French sources from Common Corpus, not only the cultural heritage part but also some of the administrative sources (Gouvert et al., 2025). In addition, NeKo, a generalist language correction model trained by Nvidia, used the OCR-corrected part of Common Corpus both for training and evaluation (Lin et al., 2024). Anthropic also used Common Corpus in their experiments for feature visualization (Ameisen et al., 2025).
Common Corpus is also being used to create new datasets. For instance, YouTube Commons was used to create two major multimodal datasets: FineVideo (43000 YouTube videos) [^17] https://huggingface.co/datasets/HuggingFaceFV/finevideo and Mosel (1 million audio hours, half from YouTube Commons) (Gaido et al., 2024), which was used to train the FAMA series of speech foundational models (Papi et al., 2025).
6 Conclusion
Common Corpus is a significant contribution to the open LLM research community. Since the release of the first version, the data from Common Corpus has been used in numerous LLM projects by several European labs and industry leaders like Anthropic. Through the release of Common Corpus and this paper with thorough documentation of data collection and curation, we show that LLM development is possible while strictly adhering to the regulatory norms. We hope that Common Corpus will grow as a critical infrastructure for open science LLM research and development and inspire future initiatives in the open.
Limitations
Common Corpus is far from collecting the whole range of available open data, which we call the open data paradox. Continuing the work, we are currently collecting more data in European languages as part of the LLMs4EU project [^18] https://www.openaire.eu/llms4eu. In addition to this, even though the language diversity of the data is large, the coverage for low-resource languages is still shallow. We are developing in this direction, planning for enlargement of the corpora in African languages (e.g., Swahili, Wolof, Bambara). Furthermore, the amount of collected data (2 trillion tokens) is suitable for pre-training of models of limited size, as in the Pleias model family (see Section 5). Larger models, however will require significantly larger amounts of data. Therefore, future collection of permissible data is highly encouraged by this work.
In addition, Common Corpus naturally does not contain data for instruction-tuning and any forms of specialized tasks. Therefore, it is not directly suitable for task-specific fine-tuning. However, due to the multilingual, temporal, and semantic diversity of data, Common Corpus opens the opportunities for the creation of ethical fine-tuning datasets, as we did for the RAG models (see Section 5).
In Section 4, we described the tools we used for the data curation, filtering, and editing. Even though we used these methods responsibly and mitigated many issues overlooked by the counterparts (e.g., with toxicity detection and PII removal), none of the curation methods could naturally facilitate a hundred-percent accuracy. However, some issues, like OCR errors, present considerable challenges to the models and might even account for better handling of typos in the future. We would also like to mention that each data object is accompanied by sufficient metadata, and, if desired, LLM practitioners are free to filter out collections that might contain potential issues (as described in Section 4).
Acknowledgments and Disclosure of Funding
Common Corpus is part of the Current AI initiative to accelerate the development of the building blocks of open AI — notably data — that serves and is shaped by the public interest. It was built up with the support and concerted efforts of AI Alliance, the state start-up LANGU:IA (start-up d’Etat), supported by the French Ministry of Culture and DINUM, as part of the prefiguration of the service offering of the Alliance for Language technologies EDIC (ALT-EDIC). The creation of the OpenCulture datasets was made possible by the support of LANGU:IA. The Wikidata and Wikipedia datasets were made in partnership with Wikimedia Enterprise and Wikidata/Wikimedia Germany. The corpus was stored and processed with the generous support of Genci (Jean-Zay, Idris, Eviden), Scaleway, and Tracto AI. The collection of the corpus has been largely facilitated thanks to major organizations committed to an open science approach for AI, namely AI Alliance, Mozilla, HuggingFace, Occiglot, and Eleuther AI.
Appendix A Development
Common Corpus was first conceptualized in May 2023 as a French corpus for the French public services. The initial aim was to create a dataset of 100-140B tokens and use that data to train a reproducible small language model, modeled on the Pythia suite (Biderman et al., 2023). The initial nucleus of Common Corpus was French-PD a large collection of 85B words from monographs and newspapers in the public domain [^19] Released as https://huggingface.co/datasets/PleIAs/French-PD-Books and https://huggingface.co/datasets/PleIAs/French-PD-Newspapers..
A collection of various open datasets was first published under the name Common Corpus in March 2024202020Release blogpost: https://huggingface.co/blog/Pclanglais/common-corpus. At the time, it was a multilingual expansion of French-PD, retaining the initial focus on cultural heritage data but containing now about 500B words in the public domain from a variety of open collections in English, German, French, Italian, Spanish and other European languages. Throughout 2024, additional official releases gradually expanded this initial pretraining commons beyond the public domain, including Youtube Commons (in April 2024) and Finance Commons (in July 2024).
In November 2024, Common Corpus became an unified corpus, through the convergence of newly collected datasets by PleIAs and pre-existing ones (especially Wikimedia Projects). It contained two trillon tokens with the current division in five collections [^21] Release blogpost: https://huggingface.co/blog/Pclanglais/two-trillion-tokens-open. The current version of Common Corpus, the one we describe in this paper, was released in February 2025 for the Paris AI Summit. Common Corpus will remain an evolving resource reflecting the need for a continuous open infrastructure rather than a fixed dataset, with HuggingFace ensuring a proper archiving of past versions for reproducibility.
A primary driver of Common Corpus has been what we term the open data paradox: most of the content available under free licenses or uncopyrighted is not integrated into the existing pretraining datasets. Since GPT-3, the pretraining data infrastructure has been extensively reliant on web archiving and the constraints that stem from it. which include the lack of available tools for PDF hosting and processing, or for diverse language detection.
The open data paradox had the effect of making the open pretraining ecosystem viable in the first place. In particular, the gradual filtering of web archives according to more and more ethical rules leaves only a tiny residual amount under open licenses (less than 2% of the original set).
Appendix B Provenance
B.1 Open Govenment
In this section, we describe the provenance and present token counts and main languages for the two sub-collections of Open Government: Finance Commons and Legal Commons.
B.1.1 Finance Commons
Dataset Main Languages Documents Tokens
SEC English 1,085,113 09,653,919,837
WTO English, Spanish, French, and small partitions of others 0,772,508 02,835,007,015
AMF French, English 0,595,397 09,823,755,281
TED EU Tenders German, French, Polish, Spanish, Dutch, Czech, Romanian, English, Swedish, Italian, Bulgarian, Finnish, Latvian, Danish, Lithuanian, Croatian, Estonian, Hungarian, Portuguese, Slovenian, Slovak, Greek, Irish 0,137,837 00,650,396,761
GATT Library English, French, Spanish, Catalan, Portuguese, German 0,067,596 00,224,526,628
Table 4: Finance Commons sources distribution with languages.
The datasets that make up Finance Commons [^22] https://huggingface.co/collections/PleIAs/finance-commons-66925e1095c7fa6e6828e26c are presented in Table 4. Here, we also present the provenance details for each of the parts of Finance Commons:
  • Securities and Exchange Commission (SEC). This dataset comprises the SEC annual reports (Form 10-K) for the years 1993 to 2024. Entries up to 2020 were compiled by Loukas et al. (2021). We added the reports from 2021-2024, which come from the EDGAR database [^23] https://www.sec.gov/search-filings/edgar-search-assistance/accessing-edgar-data, compiled using the EDGAR-Crawler toolkit [^24] https://github.com/nlpaueb/edgar-crawler.
  • World Trade Organization (WTO). This dataset comprises documents from WTO’s official Documents Online platform [^25] https://docs.wto.org/dol2fe/Pages/FE_Search/FE_S_S005.aspx. The documents cover the years 1995 to 2024. Documents are available in three official languages: English, French, and Spanish. Some documents are available in other languages, e.g., Chinese, Korean, Arabic, German, and Portuguese. Also released separately as WTO-PDF [^26] https://huggingface.co/datasets/PleIAs/WTO-PDF.
  • French Authority for Financial Market (AMF). This is a dataset of documents from the French Authority for Financial Market, or the Autorité des marchés financiers [^27] https://www.amf-france.org/en/news-publications/publications/open-data (AMF), which is an independent public authority that regulates the French market. The documents are primarily in French. Also released separately as AMF-PDF [^28] https://huggingface.co/datasets/PleIAs/AMF-PDF.
  • Tenders Electronic Daily (TED) EU Tenders. This dataset is a collection of procurement notices published by the EU. The documents are published in the online version of the “Supplement to the Official Journal” of the EU [^29] https://ted.europa.eu/en/, dedicated to European public procurement. The documents are mostly in German, with French, Polish, and Spanish making up relatively large portions of the remaining documents. There are also small portions of other languages (see details in Table 4).
  • General Agreement on Tariffs and Trade (GATT) Library. This dataset comprises documents from GATT, which was an organization that promoted international commerce and the reduction of trade barriers among member states. Public documents were made available by the General Council of the WTO in 2006303030https://www.wto.org/english/docs_e/gattdocs_e.htm. The documents span from January 1, 1946, to September 6, 1996. Most of the documents are in English, but there are also documents in French, Spanish, and other languages.
B.1.2 Legal Commons
Dataset Languages Tokens
Caselaw Access Project English 013,821,842,995
Court Listener English 022,625,121,735
EUR-lex Bulgarian, Croatian, Czech, Danish, Dutch, English, Estonian, Finnish, French, German, Greek, Hungarian, Irish, Italian, Latvian, Lithuanian, Maltese, Polish, Portuguese, Romanian, Slovak, Slovenian, Spanish, Swedish 065,044,763,781
Eurovoc English, German, French, Croatian, Italian, Lithuanian, Portuguese, Finnish, Danish, Bulgarian, Dutch, Polish, Greek, Swedish, Hungarian, Czech, Spanish, Maltese, Latvian, Slovak, Slovenian, Romanian, Estonian, Arabic, Tigrinya, Farsi, Russian, Urdu, Serbian, Albanian, Kurdish, Pushto, Irish, Norwegian, Icelandic, Dari, Armenian, Japanese. 031,648,136,898
French open data French 024,597,392,089
USPTO English 200,509,900,178
UN Digital Library Arabic, Chinese, English, French, Russian, Spanish 001,781,037,875
European Open Data EU languages 007,098,502,579
OECD English, French 000,584,969,458
Table 5: Legal Commons sources distribution with languages.
Here, we present the provenance details for each of the parts of Legal Commons:
B.2 Open Culture
Corpus Language Domain Tokens
English PD English Books and Newspapers 174.2B
US PD Books English Books 082.2B
French PD Books French Books 024.0B
French PD Newspapers French Newspapers 110.8B
French PD Diverse French Books and Newspapers 069.6B
LoC Books English Books 010.6B
US PD Newspapers English Newspapers 199.3B
New Zealand PD Newspapers English, Māori Newspapers 012.6B
Europeana Newspapers Multilingual Newspapers 021.0B
German PD Newspapers German Newspapers 018.4B
German PD German Books 058.0B
Portuguese PD Portuguese Books and Newspapers 002.6B
Spanish PD Newspapers Spanish Newspapers 008.0B
Spanish PD Books Spanish Books 015.4B
Italian PD Italian Books 018.2B
Dutch PD Dutch Books and Newspapers 002.7B
BnL Newspapers German, French, Luxembourgish Newspapers 000.3B
Danish PD Danish Books and Newspapers 000.5B
Serbian PD Serbian Books and Newspapers 000.3B
Czech PD Czech Books and Newspapers 000.7B
Greek PD Greek Books and Newspapers 004.2B
Multilingual PD Multilingual Books and Newspapers 008.4B
Polish PD Polish Books and Newspapers 005.9B
Latin PD Latin Books 027.2B
Russian PD Russian Books 001.9B
Arabic PD Arabic Books 000.3B
Table 6: Subsets of Open Culture with language coverage, domains, and token counts.
Large portion of data in Open Culture part of the Common Corpus was built on top of the following collection-as-data initiatives:
  • Chronicle America: about 100B words (150B tokens) of digitized US newspapers by the Library of Congress, made available as a raw text file.
  • Europeana: about 21B tokens of digitized European newspapers through large-scale cross-national contributions and new digitizations.
  • Gallica: about 85B words of digitized French newspapers and monographs made available on the open data portal of the French digitized library through entire dumps or API access [^46] https://api.bnf.fr.
  • Biblioteca: about 15B words of digitized Spanish newspapers and monographs.
Combined with the other retrieved data, the collections were dispatched into smaller individual subsets, which were also separately released as parts of the Open Culture collection (Table 6). The Open Culture data in Common Corpus have been post-processed and filtered, as described below, which results in a slightly different final word and token count:
  • French PD. This corpus is based on the training corpus for gallicagram [^47] https://shiny.ens-paris-saclay.fr/app/gallicagram. It comprises 289,000 books from the French National Library (Gallica). This initial aggregation was made possible thanks to the open data program of the French National Library and the consolidation of public domain status for cultural heritage works in the EU following the 2019 Copyright Directive (Art. 14).
  • French PD Newspapers. This dataset was also based on the Gallicagram corpus. It comprises nearly three million unique newspaper and periodical editions from the French National Library (Gallica).
  • LoC Books. This dataset comprises 140,000 English books, digitized by the Library of Congress. The books come from the Selected Digitized Books Collection [^48] https://www.loc.gov/collections/selected-digitized-books/about-this-collection/. The dataset was curated by using the Library of Congress JSON API. This dataset contains only the books in the English collection. The dataset was compiled by Sebastian Majstorovic.
  • US PD Newspapers. This dataset comprises 21 million digitized newspapers from Chronicling America [^49] https://chroniclingamerica.loc.gov/. The newspapers were digitized by the Library of Congress. The dataset can be fully explored through an original corpus map created by Nomic AI [^50] https://atlas.nomic.ai/data/aaron/pdnews-21286k-tr2k-addmeta/map. The dataset is mostly in English, but it also contains articles in other languages, mostly German and Spanish. The articles were published between the years 1690 and 1963.
  • New Zealand PD Newspapers. This dataset comprises historic newspapers from New Zealand and the Pacific from the 19th and 20th centuries. The data were made available by the National Library of New Zealand as part of Papers Past [^51] https://paperspast.natlib.govt.nz/newspapers. The articles are primarily in English, but include some articles in te reo Māori.
  • Europeana Newspapers. This dataset contains over 1,000 digitized newspapers from 23 libraries around Europe. It contains articles in at least 40 languages, and its articles were published between 1618 and 1990 (Neudecker, 2016). The original sources are available via Europeana, and were made available by Big Science [^52] https://huggingface.co/datasets/biglam/europeana_newspapers.
  • German PD Newspapers. This dataset contains articles from 4,299,653 issues from over 1900 different newspapers. The articles come from the German Digital Library, hosted by Deutsches Zeitungsportal [^53] https://www.deutsche-digitale-bibliothek.de/newspaper. The articles were originally published between 1794 and 1957. This dataset was curated and first made available by Sebastian Majstorovic [^54] https://huggingface.co/datasets/storytracer/German-PD-Newspapers.
  • German PD. This dataset contains texts from various sources, including the Mannheim Corpus of Historical Newspapers and Magazines [^55] https://repos.ids-mannheim.de/fedora/objects/clarin-ids:mkhz1.00000/datastreams/CMDI/content (Mannheimer Korpus Historischer Zeitungen und Zeitschriften). This dataset is made up of 21 German newspapers and magazines. The texts were originally published between 1737 and 1905. The corpus was originally digitized between 2009 and 2011. The corpus was made available by the Institut für Deutsche Sprache in 2013.
  • Spanish PD Books. This dataset contains 302,640 individual texts from various sources, including the leading cultural heritage institution Biblioteca Digital Hispánica [^56] https://www.bne.es/fr/catalogues/biblioteca-digital-hispanica (BDH). To ensure that these texts are in the public domain, we have retained exclusively titles published prior to 1884.
  • Dutch PD. This dataset contains approximately 176,000 books and 540,000 periodicals, which come from various sources including Delpher [^57] https://www.digitisednewspapers.net/histories/delpher/. Delpher is a repository of digitized printed material from the Netherlands, which is maintained by the Koninklijke Bibliotheek, the national library of the Netherlands. To ensure that these texts are in the public domain, we have retained exclusively titles published prior to 1884.
  • BnL Newspapers. This dataset contains 630,709 articles from 21 different newspaper titles and 24,415 unique issues. The articles were digitized by the National Library of Luxembourg (BnL) as part of their Open Data Initiative [^58] https://data.bnl.lu/. OCR was done using Nautilus-OCR [^59] https://github.com/natliblux/nautilusocr. The articles are in German, French, and Luxembourgish. The newspapers were originally published between 1841 and 1879. The dataset was published and made accessible by BigScience.
  • The rest of the datasets, including French PD Diverse, Portuguese PD, Italian PD, Polish PD, Danish PD, Swedish PD, Serbian PD, Czech PD, and Multilingual PD, come from various sources, including several European national libraries and cultural heritage institutions. To ensure that these texts are in the public domain, we have retained exclusively titles published prior to 1884.
B.3 Open Science
In Table 7, we present the total token counts per collection inside of the Open Science part of Common Corpus.
Dataset Tokens
OpenAlex 191,616,437,384
Open Science Pile 011,096,766,324
Open Science French 046,961,690,792
Open Science Spanish 016,523,491,767
Open Science German 007,806,446,050
ArXiv 007,188,731,472
Total 281,193,563,789
Table 7: Token count by dataset Open Science.
B.4 Open Code
Language Tokens
Java 35,697,451,454
JavaScript 28,894,772,110
Python 26,681,331,771
C++ 25,481,950,314
C 23,277,000,113
PHP 23,077,121,733
C# 16,806,995,110
Go 11,200,587,099
Rust 03,888,428,173
Ruby 03,718,918,983
Table 8: Token counts by programming language or framework.
Table 8 shows the number of tokens for the top ten coding languages and frameworks in Open Code.
Appendix C Open Culture Verification
Here, we describe the rights verification process that we applied for cultural data objects:
  • Author life + 70 years for all non-US authors. Among most signatories of the Berne Convention for the Protection of Literary and Artistic Works [^60] https://www.wipo.int/treaties/en/ip/berne/, this is the most common approach to determining documents in the public domain. This approach requires not only identifying the author but also their date of death. On top of the information already made available by cultural heritage institutions, we also implemented an internal data reconciliation pipeline based on the complete dump of Wikidata.
  • All publications after 1884. In cases where the author could not be identified or for collective works like newspapers, we applied a “universal” public domain rule based on 70 years prior to the current term of the author’s life + 70 years. Simplified rules like these are commonly applied in cultural heritage projects, especially for the release of newspaper collections.
  • Publication + 95 years for US authors. This is the copyright-based approach currently in place in the US. For an international project, this will only affect US-born authors. Due to a lack of further legal expertise, we did not attempt to include works whose copyright might not have been renewed.
  • No digitization rights. Following on the 2019 Copyright Directive (Art. 14) and common practice among GLAM reusers like Wikimedia Commons, we consider that the simple act of digitization does not provide any additional rights.
Appendix D Cleaning and Curation
D.1 Text Segmentation
Here is an example input text for the Segmentext model:
In this respect, the insurance business investment portfolio can be considered conservatively managed as it is largely composed of corporate, sovereign, and supranational bonds, term loans as well as demand deposits. Following the previous year, the group continued to diversify its holdings into investment-grade corporate bonds. It should be noted that bonds and term loans are held to maturity in accordance with the group’s business model policy of "inflows".
Technical liabilities on insurance contracts.
The guarantees offered cover death, disability, redundancy, and unemployment as part of a loan protection insurance policy. These types of risk are controlled through the use of appropriate mortality tables, statistical checks on loss ratios for the population groups insured, and through the insurance program.
Liability adequacy test.
A goodness-of-fit test aimed at ensuring that insurance liabilities are adequate with respect to current statements of future cash flows generated by the insurance contracts is performed at each statement of account. Future cash flows resulting from the contracts take into account the guarantees and options inherent therein. In the event of inadequacy, the potential losses are fully recognized in the income statement. The modeling of future cash flows in the insurance liability adequacy test are based on the following assumptions: At the end of 2022, this liability adequacy test did not reveal any anomalies.
Income statement.
The income and expenses recognized for the insurance contracts issued by the group appear in the income statement in "Net income of other activities" and "Net expense of other activities".
Risk management.
The group adopts a "prudent approach" to its management of the risks to which it could be exposed through its insurance activities. Risk of counterparty. As stated above, insurance companies only invest in assets (bank deposits, sovereign bonds, supranational agencies, or corporate bonds).
Example output:
D.2 OCR Error Detection
OCRoscope.
To illustrate this approach, this long text is correctly identified as French with >99% confidence by cld2, as despite the many mistakes, there are enough non-ambiguous French words:
NOUVELLES POLI TI QÛ E S. Suede. Stockholm , le 2 5 décembre 1792. Le général Toll ira à Varsovie en quarté d’envoyé de la Suede auprès du roi et de la république ; A 1 même rey.u l’ordre de s’y rendra incessamment. 11 paraît que k Uc-régeik a des craintes ; il a fait venir chez lji les membres c Ij“‘ tribunal 4e la cour , et leur a rtmis son lesfca n at. La fermentation qu’a causée 1 ,’ari r?tavh n k M p v riote Thorild tî’est pas appaisée y le luigage qv’il a yailé an duc-régent a été bien entendu par le peu) k y ir M» (U i n’entendrait pas l’apostrdphe suivante ? ttRxc3xa7nd >la libuk à r otre raison , et ne et nous force pas de i’ache’ef r i te n :e sang,.
Le duc a fait x,épa4idre sur-le-champ une fjtbprijuun à te us les habitans di\$ Toyaume , pour les detourntr de mr laisser sé luire par de fa,ux bruits et des jugemens pe rver\$ , e i en même temps l’ordre a. été donné à la garnison de charger et de se tenir prête à marcher.
(Mercure Français, 1793, January 25th)
Yet one short n-gram ("n k M p v riote Thorild") is classified as unknown by cld2.
OCRerrcr.
The following is a low-error example sentence taken from Common Corpus:
They did not approach cer, but turned away and passed irom her presence, filled with sorrow and moved with sympathy, which her intense emotions seemed to communicate to even these thoughtless young men of the tho plains.
And the OCRerrcr detection (with formatting for clarity):
They did not approach <<<er>>>cer,<<</er>>> but turned away and passed <<<er>>>irom<<</er>>> her presence, filled with sorrow and moved with sympathy, which her intense emotions seemed to communicate to even these thoughtless young men of the <<<er>>>tho<<</er>>> plains.
D.3 OCR Correction
Here is an example of text containing various OCR errors:
Theguaran tees offered cover death,disability,r e dundancy andunem ployment aspartof aloanprotect ion insurance policy. These types o f risk are controlled throu ghthe use o f app ropriate morta litytables,statistica lchecksonloss rat ios for thepopulation groups insure dandthrough ar e insurance program.
And here is the text corrected by our model, OCRonos:
The guarantees offered cover death, disability, redundancy, and unemployment as part of a loan protection insurance policy. These types of risk are controlled through the use of appropriate mortality tables, statistical checks on loss ratios for the population groups insured, and through the insurance program.

### v2

@@
 Abstract
-Large Language Models (LLMs) are pre-trained on large amounts of data from different sources and domains.
-These data most often contain trillions of tokens with large portions of copyrighted or proprietary content, which hinders the usage of such models under AI legislation.
-This raises the need for truly open pre-training data that is compliant with the data security regulations.
-In this paper, we introduce Common Corpus [^1] https://huggingface.co/datasets/PleIAs/common_corpus, the largest open dataset for language model pre-training.
-The data assembled in Common Corpus are either uncopyrighted or under permissible licenses and amount to about two trillion tokens.
-The dataset contains a wide variety of languages, ranging from the main European languages to low-resource ones rarely present in pre-training datasets; in addition, it includes a large portion of code data.
+Large Language Models (LLMs) are pre-trained on large data from different sources and domains.
+These datasets often contain trillions of tokens, including large portions of copyrighted or proprietary content, which raises questions about the legal use of such models.
+This underscores the need for truly open pre-training data that complies with data security regulations.
+In this paper, we introduce Common Corpus, the largest open dataset for LLM pre-training.
+The data assembled in Common Corpus are either uncopyrighted or under open licenses and amount to about two trillion tokens.
+The dataset contains a wide variety of languages, ranging from the high-resource European languages to some low-resource languages rarely represented in pre-training datasets.
+In addition, it includes a large amount of code data.
 The diversity of data sources in terms of covered domains and time periods opens up the paths for both research and entrepreneurial needs in diverse areas of knowledge.
-In this technical report, we present the detailed provenance of data assembling and the details of dataset filtering and curation.
-Being already used by such industry leaders as Anthropic and multiple LLM training projects, we believe that Common Corpus will become a critical infrastructure for open science research in LLMs.
+In this paper, we present the detailed provenance of data assembling and the details of dataset filtering and curation.
+We train two small language models on Common Corpus and find that they perform comparably to other models of their size, indicating that our dataset is suitable for multilingual pretraining.
+Common Corpus represents a key contribution to the ecosystem for open science research on Large Language Models.
+[Uncaptioned image] PleIAs/common_corpus
 1 Introduction
-Large Language Models have been defined by large training data.
-While there are several candidates for the first modern language model based on transformer architecture, including GPT-1 (Radford et al., ), ULMFIT (Howard and Ruder, 2018), or Sentence Neuron (Radford et al., 2017), it is commonly acknowledged that “large” models start with GPT-3 (Brown et al., 2020).
-Requiring a corpus of 300 billion tokens, GPT-3 introduced a standard training data pipeline shared by nearly all language models to date: large-scale processing of web datasets (45 TB of compressed source data from Common Crawl) and additional digitized sources (e.g.
-Books3).
-Until 2025, LLM training data has grown on a logarithmic curve.
-The latest generation of publicly documented language models including DeepSeek v3 (DeepSeek-AI, 2025), Gemma 3 (Google, 2025), Llama 4 (Meta AI, 2025) or Qwen 3 (Qwen, 2025) have been trained on 14-36 trillion tokens.
-Even the gaining popularity sub-category of small language models (Wang et al., 2025) relies on large amounts of training data to fit scaling laws: Qwen 3 0.6B was trained on 36 trillion tokens, about 3,000 times more data than recommended by the original Chinchilla laws (Hoffmann et al., 2022).
-As data curation also became a primary concern, the collection, maintenance, processing, and filtering of data became one of the primary cost of standard language model training.
-This does not factor in even larger hidden costs: the negative externalities affecting competing markets, the digital commons, and society at large.
+Large Language Models demand large amounts of training data.
+GPT-3 (Brown et al., 2020) is considered the first “large” language model.
+Trained on 300 billion tokens, GPT-3 introduced a standard training data pipeline shared by nearly all language models to date: large-scale processing of web datasets (45 TB of compressed source data from Common Crawl) and additional digitized sources (Books3).
+Until 2025, LLM training data sizes have grown logarithmically.
+The latest generation of publicly documented language models including DeepSeek v3 (Liu et al., 2025), Gemma 3 (Kamath et al., 2025), Llama 4 (Meta, 2025) or Qwen 3 (Yang et al., 2025) have been trained on 14-36 trillion tokens.
+Even “small language models” (Wang et al., 2025) rely on large amounts of training data to fit scaling laws: Qwen 3 0.6B was trained on 36 trillion tokens, which is 3,000 times more than Chinchilla-optimal estimates (Hoffmann et al., 2022).
+Figure 1: Proportional treemap of Common Corpus collections and their most popular languages.
+As data quality has increasingly been an area of concern, the collection, maintenance, processing, and filtering of data has become one of the main costs in language model training.
+But data curation at scale may also have hidden costs: negative externalities affecting competing markets, the digital commons, and society at large.
 While data scraped from the web is publicly available, it is not always in the public domain.
-Most web data do not have sufficient metadata to determine whether it is permissively licensed.
+Most web data does not have sufficient metadata to determine whether it is permissively licensed.
 NLP practitioners have relied on the protection of fair use, claiming that the transformative nature of the use of the data allows them to use this data to train language models.
 There are increasingly more legal challenges to the use of this data.
-The New York Times sued Open AI for copyright infringement, alleging that Open AI trained their models on their articles (Roth, 2023, Pope, 2024).
+The New York Times sued OpenAI for copyright infringement, alleging that OpenAI trained their models on NYT articles  (Roth, 2023; Pope, 2024).
 Due to concerns about indirect commercial exploitation, many rightholders have implemented either hard technical measures or legal provisions against model training.
 In 2024, it was estimated that for Terms of Service crawling restrictions, a full 45% of C4 is now restricted (Longpre et al., 2024b) and 5% is fully blocked for scraping with a disproportionate impact over quality sources (Longpre et al., 2024b).
-Restrictions not only affect language model pre-training but also the quality of search engine indexation and a variety of research projects analyzing and collecting content at scale.
+Restrictions not only affect LLM pre-training but also the quality of search engine indexation and a variety of research projects analyzing and collecting content at scale.
 Even projects dedicated to knowledge access have faced significant pressure from AI crawlers and implemented protections that negatively impact access and user experience.
-These legal uncertainties have significantly impeded the development of open science research on LLMs.
+Legal uncertainties have significantly impeded the development of open science research on LLMs.
 Previously reproducible research artifacts have been removed or taken down, impacting pre-training data, continuous pre-trained models, and evaluation datasets.
-Books3, which has been used in datasets like the Pile (Gao et al., 2020) and by multiple AI labs, including Meta and Microsoft, faced legal challenges (Brittain, 2023), and the original dataset was ultimately removed (Van der Sar, 2023).
-The LAION dataset was demonstrated to contain CSAM (Birhane et al., 2021, Thiel, 2023), and then was taken down (LAION, 2023).
-Later, it was re-released once suspected CSAM was removed (LAION, 2024).
-The Dutch-focused model GEITje was taken down (Rijgersberg, 2025) due to complaints about it having been trained on the Dutch Gigacorpus, in order to avoid legal disputes.
+Books3, which has been used to train numerous models, faced legal challenges (Brittain, 2023).
+The original dataset was ultimately removed due to a DMCA take-down (Van der Sar, 2023).
+The LAION dataset was demonstrated to contain CSAM (Birhane et al., 2021; Thiel, 2023), and taken down (LAION, 2023), and then re-released once suspected CSAM was removed (LAION, 2024).
+The Dutch model GEITje was taken down (Rijgersberg, 2025), due to complaints about training on the Dutch Gigacorpus, in order to avoid legal disputes.
 Finally, the widely used benchmark, the Mathematics Aptitude Test of Heuristics (MATH) dataset (Hendrycks et al., 2021), was removed from Hugging Face via a DMCA takedown.
@@
 These takedowns and legal challenges also represent a sizeable loss of investment for developers, who are often independent or small research organizations.
-In part, as a reaction to the use of publicly available but not permissively licensed data, web text is becoming increasingly more difficult to acquire and use.
+In part as a reaction to the use of publicly available but not permissively licensed data, web text is also becoming harder to acquire and use.
 In an analysis of popular datasets such as C4 (Raffel et al., 2020), RefinedWeb (Penedo et al., 2023), and Dolma (Soldaini et al., 2024), Longpre et al.
-(2024c) found that just in the last year, 5% of all tokens in C4 now have restricted use, with a disproportionate number of those tokens coming from the best-maintained, most critical sources.
+(2024b) found that just in the last year, 5% of all tokens in C4 now have restricted use, with a disproportionate number of those tokens coming from the best-maintained, most critical sources.
 This is largely due to changes in content owners’ and hosts’ preferences, which are changing to no longer allow scraping, especially for the purposes of training AI models.
-Since 2024, several initiatives have emerged to collect open data with clear licensing and in the English language.
-This includes: C4C, Open License Corpus, a 228 billion token corpus from a mix of public domain texts and open source code under free licenses not requiring attribution (Min et al., 2024), KL3M, a 1.2 trillion tokens corpus of administrative texts and structured data mostly from the US federal public domain (Bommarito et al., 2025), Common Pile, an ongoing data collection of 1 trillion tokens from a variety of recent sources, including a filtered version of Common Crawl data (Creative Commons Common Crawl) (Kandpal and Raffel, 2025).
-These projects are all monolingual, restricting in effect the reach of open-data language models to the English-speaking audience.
-In contrast, the most ambitious multilingual collection of permissive content pre-dates Large Language Models: C4C (2016), containing 12 million web pages in more than 50 languages filtered by Creative Commons Licenses (Habernal et al., 2016).
-In this paper, we present Common Corpus, which has been an ongoing effort since 2024 (see Appendix A for the development history).
-It has grown to become the largest open dataset for pre-training Large Language Models at about 2 trillion tokens and the only one in its size range available in multiple languages.
-Thus, Common Corpus is a significant contribution to the AI Commons.
-Through the release of this dataset, we show that open LLM research and development is possible while meeting legal and regulatory requirements—even in the EU, where AI regulations are currently the strictest.
-Here, we detail the composition of Common Corpus (Section 3), as well as the data provenance, license verification, and further processing (Section 4).
-Next, we describe the early impact of Common Corpus on the AI community (Section 5) and highlight the limitations of the work that has been done at the time of publication.
-Despite its size, Common Corpus is still far from covering the entire range of available resources for pre-training commons: we attribute this discrepancy to an open data paradox as major sources of open content are paradoxically little visible online and even more so in the leading sources currently used for pre-training.
-By describing the unique challenges coming with the aggregation of large open source, we aim to inspire further responsible data collection initiatives.
-Figure 1: A schematic world map of languages in Common Corpus with a log-scaled distribution of document counts.
-For each language, we chose a city that is located in the region where this language is most specific to.
-To avoid outliers, we show only languages with 1000+ documents.
+Since 2024, several initiatives have emerged to collect open data in English with clear licensing.
+This includes: C4C (Habernal et al., 2016); Open License Corpus (Min et al., 2024), a 228 billion token corpus from a mix of public domain texts and open source code under free licenses; KL3M (Bommarito et al., 2025), a 1.2 trillion tokens corpus of administrative texts and structured data mostly from the US federal public domain; Common Pile (Kandpal et al., 2025), a data collection of 1 trillion tokens from a variety of recent sources, including a filtered common crawl (Creative Commons Common Crawl).
+All these projects are monolingual, restricting in effect the reach of language models to the English-speaking audience.
+In contrast, the most ambitious multilingual collection of permissive content predates Large Language Models: C4C contains 12 million web pages in more than 50 languages filtered by Creative Commons Licenses (Habernal et al., 2016).
+Common Corpus has grown to become the largest fully open pre-training dataset at about 2 trillion tokens and the only one in its size range having high multilingual diversity.
+Through this release, we show that open LLM research and development is possible while meeting legal and regulatory requirements — in compliance with even the strictest AI regulations, such as in the European Union.
+In this paper, we detail the composition of Common Corpus and the process of data collection and curation, and license clearing.
+Despite its size, Common Corpus is still far from covering the entire range of available resources.
+We attribute this discrepancy to an open data paradox as major sources of open content are paradoxically little visible online and even more so in the leading pre-training sources.
+By describing the unique challenges coming with the aggregation of large open source, we aim to inspire further initiatives.
+We also train two small language models on our dataset and find that it offers comparable performance to existing multilingual models.
 2 About Common Corpus
@@
 Not only is the data available, but we also provide essential details about the data provenance, data processing, and important information about the contents of each dataset.
-Following the Open Source Initiative’s definition of open-source AI, our data is open in terms of openness of use, meaning that use is permitted for “any purpose and without having to ask for permission” (Open Source Initiative, 2024).
-To achieve this, models must be trained on datasets that are free from copyright or other relevant legal limitations.
+The Open Source Initiative has also defined open-source AI in terms of openness of use, where open means that use is permitted for “any purpose and without having to ask for permission” (Open Source Initiative, 2024).
+To achieve this, models must be trained on datasets that are free from copyright or other legal limitations.
 This is currently a limitation of existing open datasets for training LLMs.
+Figure 2: A schematic world map of languages in Common Corpus with a log-scaled distribution of document counts.
+For each language, we chose a city that is located in the region where this language is most specific to.
+To avoid outliers, we show only languages with 10,000+ documents.
 Common Corpus, therefore, provides valuable training tokens that will not be subject to the same restrictions.
 Additionally, the data in Common Corpus are different from other corpora, primarily composed of web text.
-Common Corpus contains multilingual data in a variety of high- and low-resource languages (see Figure 1 for language distribution), covering diverse genres, time periods, and domains (in Section 3, we detail each part of the dataset).
-Therefore, Common Corpus contributes to data diversity in the open pre-training data ecosystem.
+It contains multilingual data in a variety of high- and low-resource languages (see Figure 2 for language distribution), covering diverse genres, time periods, and domains (in Section 4, we detail each part of the dataset).
+Therefore, it contributes to data diversity in the open pre-training data ecosystem.
 This is important for developing powerful and generalizable model performance.
 Common Corpus can be used on its own or in conjunction with existing open datasets, according to one’s needs and the desired use case of a language model.
-(a) A timeline of the main collections with their numbers of documents in the Common Corpus.
-(b) A two-component t-SNE visualization of a subset of the Common Corpus.
-Figure 2: Temporal and semantic overview of the Common Corpus collections.
-Dataset Documents Words Tokens
-Open Government 074,727,536 0,257,233,670,261 0,406,581,454,455
-Open Culture 093,156,602 0,549,608,763,966 0,885,982,490,090
-Open Science 019,220,942 0,147,305,783,453 0,281,193,563,789
-Open Code 202,765,051 0,077,669,169,092 0,283,227,402,898
-Open Web 096,165,348 0,033,208,509,065 0,073,217,485,489
-Open Semantic 030,072,707 0,023,284,201,782 0,067,958,671,827
-Other 000,925,462 0,000,328,160,421 0,000,486,099,734
-Total 517,033,648 1,088,638,258,040 1,998,647,168,282
-Table 1: Dataset composition of Common Corpus. For each collection, we report the total number of documents, words (whitespace-separated), and tokens.
-Common Corpus was developed with consideration for ongoing conversations about best practices for open-source LLM development (The AI Alliance, 2024, Longpre et al., 2024a, Duprieu and Berkouk, 2024, Baack et al., 2025).
-As PleIAs was one of the contributors at the Dataset Convening, we highlight our adherence to the best practices that were suggested by Baack et al.
+Common Corpus was developed with consideration for developing best practices for open-source LLM development (The AI Alliance, 2024; Longpre et al., 2024a; Duprieu and Berkouk, 2024; Baack et al., 2025).
+We highlight our adherence to those suggested by Baack et al.
 (2025):
   • Provide useful documentation.
-We provide information about dataset provenance and processing (Sections 3 and 4) and share key statistics to help potential users understand the applications of the dataset.
-This is crucial, as dataset documentation improves reproducibility, helps prevent misuse, and aids downstream users to best utilize the dataset (Longpre et al., 2024a).
+We provide information about dataset provenance and processing (Sections 4 and 5) and share key statistics to help potential users understand the applications of the dataset.
+Dataset documentation improves reproducibility, helps prevent misuse, and aids downstream users to best utilize the dataset (Longpre et al., 2024a).
   • Follow and record preference signals. In the metadata, we include the source URL and license information for the vast majority of the corpus.
@@
 In particular, for the dataset in the public domain, we engage in extensive OCR correction and toxicity filtering in order to bring datasets up to standard (Section 5).
-  • Do not capture highly sensitive data. We remove personally identifiable information from our datasets (Section 4.4).
+  • Do not capture highly sensitive data. We remove personally identifiable information from our datasets (Section 4).
 Common Corpus aims to support the pre-training of fully open and auditable LLMs by making it legal to release the source even without the provision of fair use.
@@
 We argue this is the first point in time where there has been sufficient knowledge and infrastructure to collect and clean a dataset on this scale, which meets the legal and ethical criteria we have outlined.
-2.1 Composition
-Common Corpus is available on HuggingFace as an aggregation of 10,000 parquet files and is composed of six collections: Open Government, Open Culture, Open Science, Open Web, Open Code, and Open Semantic.
-In total, the number of tokens in Common Corpus is 1,998,647,168,282.
-The token counts [^2] We report token counts in terms of the Pleias base tokenizer https://huggingface.co/PleIAs/Pleias-350m-Preview trained on a subsample of Common Corpus.
-in each collection are listed in Table 1.
-We visualize the timeline of the collected documents and embeddings of a subsample in Figure 2.
-Each collection is composed of multiple datasets, for which we provide details about provenance and other key information in the corresponding subsections.
-Each data object contains a license, language(s), a collection/domain of specialization, and other metadata, allowing one to filter out a desired subset.
-Language Tokens
-English 968,757,721,747
-French 275,358,437,630
-German 112,127,458,251
-Spanish 046,514,142,421
-Latin 036,031,591,540
-Italian 024,681,637,575
-Polish 012,146,688,669
-Greek 011,376,498,056
-Portuguese 010,262,747,943
-Russian 009,439,453,633
-Table 2: Token counts for the ten most represented languages in Common Corpus.
-License type Tokens
-Public Domain 1,138,508,375,958
-$C$-By 0,287,749,264,457
-MIT 0,142,694,227,607
-$C$-By-SA 0,074,768,060,836
-Apache-2.0 0,068,750,977,037
-BSD-3-Clause 0,018,483,944,333
-Open license 0,010,432,513,767
-BSD-2-Clause 0,005,497,145,480
-$C$-BY-4.0 0,002,110,966,243
-CC0-1.0 0,001,877,206,195
-Table 3: Token counts for the ten most common licenses in Common Corpus.
-The majority of the data in Common Corpus is in the public domain (see Table 3).
-The license for each document is provided in the metadata, so the dataset can be easily filtered by license as needed for a particular use case.
-Common Corpus is multilingual (see Figure 1 and Table 3 for language coverage).
-Unlike many pretraining mixes, which are coposed mostly of English and code data, our dataset has strong representation for several languages, with the top nine languages constitute at least 10B tokens each [^3] Language distribution was computed using the fastText language identification model..
-While we highlighted some of the issues faced in making open datasets above, all of these issues are exacerbated for languages other than English.
+Common Corpus is composed of six collections: Open Government, Open Culture, Open Science, Open Web, Open Code, and Open Semantic (Figure 1), and in total amounts to 1,998,647,168,282 tokens [^1] The token counts are calculated using PleIAs tokenizer trained on a representative subsample of Common Corpus, available at: [Uncaptioned image] PleIAs/Pleias-350m-Preview.
+The token counts in each collection are listed in Appendix B.
+To highlight the diversity, we visualize the timeline of the collected documents and embeddings of a subsample in Figure 3.
+Each collection comprises multiple datasets, for which we provide details on provenance and other key information in the corresponding subsections.
+Each data object contains a license, language(s), a collection or domain of specialization, and other metadata, allowing one to filter out a desired subset.
+We note that the majority of the data in Common Corpus is in the public domain (see Appendix B).
+(a) A timeline of the main collections with their numbers of documents in the Common Corpus.
+(b) A two-component t-SNE visualization of subsets of Common Corpus collections, C4, and FineWeb.
+Figure 3: Temporal and semantic overview of the Common Corpus collections.
+Common Corpus is multilingual (Figure 2 and language distribution in Appendix B), with at least 10B tokens for nine languages [^2] Language distribution was computed using the fastText language identification model..
+The issues discussed above for making datasets are compounded in languages other than English.
 Even in relatively high-resource languages like French, these problems are compounded by the fact that there is much less data available, and most tools generalize poorly to languages other than English.
@@
 (2022) showed that many multilingual datasets contain a lot of low-quality or entirely unusable data.
-Many of the datasets they analyzed contained less than 50% of usable text, with 15 sources containing no usable data at all.
+Many of the existing datasets they analyzed contained less than 50% of usable text, with 15 sources containing no usable data at all.
+3 Related Work
+Table 1 illustrates the similarities and differences between prominent open pre-training datasets and Common Corpus.
+Multiple other efforts, including C4 (Raffel et al., 2020), ROOTS (Laurençon et al., 2022), DCAD-2000 (Shen et al., 2025), and FineWeb 2 (Penedo et al., 2025), contain multilingual data from several distinct domains, but consist mostly of web crawl data with not only permissive licenses.
+Dolma (Soldaini et al., 2024) expands beyond web crawl, but is primarily comprised of English data.
+KL3M (Bommarito et al., 2025) is offering strictly permissive licensing beyond crawled data, but it is limited to administrative and legal documents in English.
+Common Pile (Kandpal et al., 2025) shares Common Corpus’s commitment to varied data with only open licensing, but is English-only.
+Common Corpus is unique in satisfying all four criteria simultaneously: multilingual and multi-domain coverage, data sources beyond web crawls, and fully open licensing.
+KL3M Dolma C4 ROOTS DCAD FineWeb 2 Common Common
+2000 Pile Corpus
+Multidomain
+Beyond Web Crawl
+Multilingual
+Open data
+Table 1: Comparison of the contemporary datasets for LLM training.
+We also stress the difference in data sources.
+An investigation of the top 1000 domains of FineWeb shows very low overlap with the Common Corpus: less than 2% of pages and 1% of domain names.
+Within overlapping collections, content can differ substantially: we focused on large PDF sets, whereas agnostic crawl pipelines mostly retrieve HTML pages (e.g., abstracts from scientific papers).
+This means that Common Corpus mostly adds new, diverse content to the open pretraining ecosystem and significantly differs from crawl corpora, as we also illustrate in Figure 3(b).
 4 Provenance
-In this section, we present the details about collections that comprise the Common Corpus, accompanied by the detailed information about the data sources, token count breakdowns, and the main included languages in Appendix B.
+In this section, we present the details about collections that comprise the Common Corpus, accompanied by the information about the data sources and the main included languages in Appendix D.
 4.1 Open Government
 Open Government is a set of financial, legal, and administrative data in the public domain.
-In total, the dataset contains more than 406B tokens and comprises two main datasets: Finance Commons and Legal Commons.
-See Appendix B.1 for detailed data composition.
+In total, the dataset contains more than 407B tokens and comprises two main datasets: Finance Commons and Legal Commons.
+See Appendix D.1 for detailed data composition.
 Finance Commons.
 This is the largest collection of financial documents in the public domain, comprising more than 14 billion words (more than 23 billion tokens).
-The documents come from a wide time range, from the mid-20th century all the way to 2024.
+The documents come from a wide time range, all the way to 2024.
 Like many of our other datasets, Finance Commons is also multilingual.
@@
 Open Culture is an aggregation of vast cultural heritage datasets containing both monographs and periodicals for over 13 languages: French, English, German, Spanish, Portuguese, Italian, Dutch, Luxembourgish, Danish, Swedish, Serbian, Czech, and Greek.
-There are also small portions of data in a wide variety of other languages, including Arabic, Bengali, Latin, Persian, Russian, Sanskrit, and Urdu.
+There are also small portions of data in other languages, such as Arabic, Bengali, Latin, Persian, Russian, Sanskrit, and Urdu.
+Composition.
 A large part of Open Culture is compiled from Collections As Data (CAD) — large dumps of texts, datasets, PDFs, and even raw XML output (METS/ALTO).
-CAD initiatives, thus, considerably simplify dataset aggregation and are a major contribution to the digital commons ecosystem.
-All other parts of Open Culture have been collected on a resource-by-resource basis using APIs and other standard retrieval methods whenever available.
-The largest extractions of this kind include Internet Archive (about 2 million monographs in multiple languages) and Delpher (50,000 Dutch monographs and periodicals filtered to match the Dutch copyright law for public domain).
-We managed to compile a large multilingual collection despite such challenges, as poor OCR quality, which we partly solved through the development of OCR correction tools (see Section 4), text segmentation issues, and sometimes irrecoverable deterioration of the original support.
-For the detailed dataset composition, refer to Appendix B.2.
-All Open Culture documents are in the public domain, which means their copyright has expired after a given term and there are no limitations on their reuse.
+CAD initiatives considerably simplify dataset aggregation and are a major contribution to the digital commons ecosystem.
+The remaining sources have been collected on a resource-by-resource basis using APIs and other standard retrieval methods whenever available.
+The largest extractions of this kind include Internet Archive (about 2 million monographs) and Delpher (50,000 Dutch monographs and periodicals filtered to match the Dutch copyright law for public domain).
+We managed to compile a large multilingual collection despite such challenges, like poor OCR quality, through the development of OCR correction tools (see Section 5), text segmentation issues, and sometimes irrecoverable deterioration of the original support.
+For the detailed dataset composition, refer to Appendix D.2.
+Licenses.
+All Open Culture documents are in the public domain, which means their copyright has expired after a given term, and there are no limitations on their reuse.
 For certain content, or in cases where we could not rely on the guarantee of established cultural heritage institutions, we implemented our own internal rights verification process.
-This process follows specific criteria, including author life and data object creation time, and takes into account that we only collected cultural heritage content from institutions based in the US or the EU (see the complete criteria list in Appendix C).
+This process follows specific criteria, including author life and data object creation time, and takes into account that we only collected cultural heritage content from institutions based in the US or the EU (see the complete criteria list in Appendix E).
+Value.
 Open Culture data is also rich from a cultural and stylistic standpoint and can be used to train multilingual language models with more diverse and creative writing styles.
@@
 In an analysis of WildChat (Zhao et al., 2024), a dataset of 1 million user interactions with ChatGPT, Longpre et al.
-(2024c) found that over 30% of user requests involved creative compositions such as fictional stories, role-play, or poetry generation.
-At the same time, creative writing is poorly represented among datasets used to train LLMs, which mainly comprise web text (Longpre et al., 2024c).
+(2024b) found that over 30% of user requests involved creative compositions such as fictional stories, role-play, or poetry generation.
+At the same time, creative writing is poorly represented among datasets used to train LLMs, which mainly comprise web text (Longpre et al., 2024b).
 Therefore, Open Culture contributes data that can be used to train models for creative writing without violating copyright law.
 In addition, as many of the Open Culture datasets are historical (coming from the 18th-19th centuries, or even earlier; see Figure 3(a)), this collection also enables the development of historical language models.
-The metadata includes document creation year, which enables researchers to develop language models with a cutoff of the training data creation date.
+The metadata includes the document creation year, which enables researchers to develop language models with a cutoff of the training data creation date.
 4.3 Open Science
 The Open Science collection includes scientific papers and other documents (theses, book reviews, clinical trials, etc).
-Following on the development of a global open access movement, these documents have been made increasingly available in open archives (preprints) or directly through open science publishers and infrastructure.
-Scientific content has become a primary focus for training data curation, due to its impact on performance on reasoning and advanced world knowledge tasks.
+Following the development of a global open access movement, these documents have been made increasingly available in open archives (preprints) or directly through open science publishers and infrastructure.
+Scientific content has become a primary focus of training data, due to its impact on reasoning capacities.
 Yet, the lack of licensing information has until now partly hindered reuse.
 The Semantic Scholar Open Research Corpus from Allen AI includes 81.1 million articles in English under an Open Data Commons Attribution License, allowing for the free reuse of the aggregated metadata while still acknowledging the remaining copyright of individual authors (Lo et al., 2020).
-The Pile incorporated dumps made available by arXiv and PubMed Central, also exclusively in English (Biderman et al., 2022).
-Finally, the BigScience project assembled several curated multilingual scientific datasets like the French HAL as part of the training data for BLOOM (Workshop et al., 2023).
-The Open Science collection was made possible largely due to the recent development of OpenAlex [^4] https://openalex.org/, the largest open catalogue of scientific documents.
+The Pile incorporated data from arXiv and PubMed Central, also exclusively in English (Biderman et al., 2022).
+Finally, the BigScience project assembled several curated multilingual scientific datasets like the French HAL as part of the training data for Bloom (Scao et al., 2023).
+The Open Science collection was made possible largely due to the recent development of OpenAlex (Priem et al., 2022), the largest open catalogue of scientific documents.
 OpenAlex maintains an expansive API search engine tracking detailed metadata for each indexed item, including the licensing, as well as a link to the original resource, which is generally in PDF format.
-We filtered OpenAlex on the three following licenses: CC-By, Public Domain/CC0, and CC-By-SA.
+We filtered OpenAlex on the following three licenses: CC-By, Public Domain/CC0, and CC-By-SA.
 The largest share of resources is available under CC-By, which is currently the recommended license by the Open Access definition.
-Open Science also includes smaller subsets, such as a direct extraction of arXiv articles available in CC-By and some European-specific resources not currently well indexed on OpenAlex (the exact distribution of token counts can be found in Appendix B.3).
+Open Science also includes smaller subsets, such as a direct extraction of arXiv articles available in CC-By and some European-specific resources not currently well indexed on OpenAlex (the exact distribution of token counts can be found in Appendix D.3).
 Due to the specificity of open scientific publishing, the Open Science collection has less linguistic diversity, with nearly 85% of documents currently available in English.
 4.4 Open Code
-The Open Code collection comprises code data under a vast variety of free licenses, which allows NLP practitioners to train models on public domain code for either coding applications or in order to improve certain model performance on natural language reasoning, world knowledge tasks, mathematics, and structured output tasks (Aryabumi et al., 2024, Petty et al., 2024, MA et al., 2024).
-The code data we use comes from the Stack v1 and v2 (Kocetkov et al., 2023, Lozhkov et al., 2024).
+The Open Code collection comprises code data under a vast variety of free licenses, which allows NLP practitioners to train models on public domain code for either coding applications or in order to improve certain model performance on natural language reasoning, world knowledge tasks, mathematics, and structured output tasks (Aryabumi et al., 2025; Petty et al., 2024; Ma et al., 2024).
+The code data we use comes from the Stack v1 and v2 (Kocetkov et al., 2023; Lozhkov et al., 2024).
 The Stack v1 contains 6.4TB of data and covers 30 programming languages, while the Stack v2 is approximately ten times bigger at 67.5TB and covers over 600 programming languages.
 All the code data is made available with a direct link to the original resource on GitHub.
-In total, Open Code contains 283,227,402,898 tokens (see most common languages in Appendix B.4).
+In total, Open Code contains 283,227,402,898 tokens (see most common languages in Appendix D.4).
 […]
@@
 Wikipedia and Wikisource.
-Wikimedia projects have always been major sources for language model training due to their reliability, extensive coverage, and textbook-like style.
+Wikimedia projects are popular sources for language model training data due to their reliability, extensive coverage, and textbook-like style.
 Despite this centrality, there is still a range of unresolved challenges with the most common versions available for training.
@@
 As part of Wikimedia Enterprise, the Wikimedia Foundation created entirely new dumps from the rendered HTML sources, which in effect ensure that they include all the text made available to readers.
-Youtube Commons.
-YouTube Commons is our collection of audio transcripts of 2,063,066 videos uploaded on YouTube under a standardized CC-By license [^5] https://huggingface.co/datasets/PleIAs/YouTube-Commons.
+Youtube Commons333[Uncaptioned image] PleIAs/YouTube-Commons.   For YouTube Commons, we collected audio transcripts of 2,063,066 videos uploaded on YouTube under a standardized CC-By license.
 StackExchange.   This is a collection of user-generated forums and Q&A made available under the CC-By-SA license. We reused the version from The Pile  (Biderman et al., 2022).
-A major objective will be the integration of web archives filtered by permissive license.
-Since 2016, several projects have attempted to reidentify Creative Commons licenses from web archives at scale including C4C (multilingual) (Habernal et al., 2016) and more recently CCCC (from Allen AI - in English) and most recently Common Crawl Creative Commons Corpus (C5; the first multilingual C4 derivative) [^6] https://huggingface.co/datasets/BramVanroy/CommonCrawl-CreativeCommons All theses projects struggled with license identification.
-While license mentions are frequently normalized with a direct link or logo to Creative Commons, there is no guarantee they really concern the entire content: “a blog page contains many photos and each photo is licensed under different CC-license type or a blog home page with many articles and each article is licensed under different CC-license type.”(Habernal et al., 2016).
+A major aim for the future work will be the integration of web archives filtered by permissive licenses.
+Since 2016, several projects have attempted to reidentify Creative Commons licenses from web archives at scale including C4C (multilingual;  Habernal et al., 2016), CCCC (English; Dodge et al., 2021), and most recently multilingual Common Crawl Creative Commons Corpus (C5; Vanroy, 2025).
+All these projects struggled with license identification.
+While license mentions are frequently normalized with a direct link or logo to Creative Commons, there is no guarantee they really concern the entire content: “a blog page contains many photos, and each photo is licensed under a different CC-license type, or a blog home page with many articles, and each article is licensed under a different CC-license type.” (Habernal et al., 2016).
 We hope this limitation could be overcome by a combination of web domain curation and fine-grained curation and annotation by a language model.
@@
 This is to date the only available textual collection of Wikidata covering the entire range of 300 languages.
-Data processing involved the translation of items and properties into formal language sequences as simple natural language sequences, without textual synthesis: “Q41309 | P:27 | Q171150” becoming “Franz Liszt country of citizenship Kingdom of Hungary”.
-Within each entry, we provide all the available translations as consecutive blocks separated by a newline, anticipating that this may contribute to language alignment.
+Data processing involved the translation of items and properties into formal language sequences as simple natural language sequences, without textual synthesis: “Q41309 — P:27 — Q171150” becoming “Franz Liszt country of citizenship Kingdom of Hungary.” Within each entry, we provide all the available translations as consecutive blocks separated by a newline, anticipating that this may contribute to language alignment.
 5 Cleaning and Curation
-4.1 Text Segmentation
-For text segmentation, we developed Segmentext [^7] https://huggingface.co/PleIAs/Segmentext (see example in Appendix D.1).
-It was trained to be resilient to broken and unstructured texts with digitization artifacts and ill-recognized layout formats.
+In order to curate our dataset, we developed a number of custom tools [^4] All described tools are publicly released: [Uncaptioned image] PleIAs/Segmentext   [Uncaptioned image] PleIAs/OCRoscope [Uncaptioned image] PleIAs/OCRerrcr   [Uncaptioned image] PleIAs/OCRonos   [Uncaptioned image] PleIAs/celadon to handle the issues unique to multilingual, historical, and OCRed data.
+We document the tools in detail in Appendix F.
+Text Segmentation.
+We developed Segmentext, a specialized language model for text segmentation (see example in Appendix F.1).
+Segmentext has been trained to be resilient to broken and unstructured texts with digitization artifacts and ill-recognized layout formats.
 Given the diversity of the training data, Segmentext should work correctly on diverse document formats in the main European languages.
-4.2 OCR Error Detection
-We developed two pipelines to determine the digitization quality of different datasets in order to determine the appropriate treatment before datasets can be used for pre-training:
-OCRoscope [^8] https://github.com/Pleias/OCRoscope.
-This is a tool based on the language identification tool cld2999https://pypi.org/project/pycld2/.
-As cld2 cannot attribute non-recognized strings to any language, OCRoscope works by measuring the number of non-recognized 7-grams.
-OCRoscope provides a standardized rate of OCR quality and rate of non-character content for at least 80 languages.
-An example of OCPoscope work is presented in Appendix D.2.
-Complete processing by OCRoscope yielded a rate of 41% non-recognized 7-grams, which results in an OCR quality rate of 59%.
-In comparison, the self-estimated rate of OCR valid words by the French National Library is significantly higher (85%) for the whole document.
-OCRerrcr [^10] https://huggingface.co/PleIAs/OCRerrcr.
-This is a DeBERTa-v2-style language model with 400M parameters.
-OCRerrcr achieves higher accuracy than OCRoscope, but is more computationally intensive.
-4.3 OCR Correction
-We developed OCRonos [^11] https://huggingface.co/PleIAs/OCRonos model based on Llama 3 8B (Grattafiori et al., 2024).
-OCRonos is versatile and supports the correction of OCR errors, wrong word cuts or merges, and overall broken text structures.
-The training data includes a highly diverse set of OCR-ed texts in multiple languages, mostly coming from uncorrected versions of Open Culture and Open Government.
-On highly deteriorated content, OCRonos can act as a synthetic rewriting tool rather than a strict correction tool.
-An example of OCRonos work is presented in Appendix D.3.
-OCRonos contributes to making challenging resources usable for LLM applications and, more broadly, search retrieval.
-It is especially fitting in situations where the original PDF sources are too damaged for correct OCR or are not possible to retrieve.
+OCR Error Detection We developed two pipelines to determine the digitization quality of different datasets in order to determine the appropriate treatment before datasets can be used for pre-training.
+The first is OCRoscope.This is a tool based on the language identification tool cld2 (Al-Rfou, 2013), which identifies documents with non-recognized 7-grams to provide a measure of OCR quality for at least 80 languages (for further details, see Appendix F.2).
+The other is OCRerrcr.
+This is a DeBERTa-v2-style language model with 400M parameters, which is trained to label OCR errors.
+This enables error rate estimation and supports OCR error correction.
+OCRoscope tends to have lower accuracy, especially for documents with fewer OCR errors.
+However, it is less computationally expensive and is faster, which allows it to be better used at larger scales.
+OCRerrcr, on the other hand, achieves higher accuracy, but is more computationally intensive.
+OCR Correction.
+We developed OCRonos model based on Llama 3 8B  (Grattafiori et al., 2024).
+It supports the correction of OCR errors, cutting or merging of the wrong word, and overall broken text structures.
+The training data includes a highly diverse set of OCR-ed texts in multiple languages, mostly from uncorrected versions of Open Culture and Open Government.
+On highly deteriorated content, OCRonos can act as a synthetic rewriting tool rather than a strict correction tool (e.g., Appendix F.4).
 OCRonos is generally faithful to the original material, provides sensible restitution of deteriorated text, and will rarely rewrite correct words.
-A common issue with OCR correction has been language switching: due to the inherent noise in the input text, an LLM will transcribe in a different language or script.
-The issue has been especially observed in smaller generalist models like GPT-3.5 or Claude-Haiku.
-OCRonos largely mitigates this issue.
-4.4 PII Removal
-Personally Identifiable Information (PII), which is any information that can be used to distinguish or trace an individual’s identity, is protected under legislation such as GDPR.
-Consequently, the new reguations put restrictions on LLM training data.
-In an analysis of large open datasets, there is a staggering amount of personal data in widely used datasets, e.g., large quantities of phone numbers in RedPajama, email addresses in S2ORC and peS2o, and IP addresses in the Stack (Elazar et al., 2024).
-To identify and replace PII, we use Microsoft’s Presidio [^12] https://microsoft.github.io/presidio/, an open-source state-of-the-art tool.
+However, when original PDF sources are too damaged for accurate OCR, unavailable, or otherwise difficult to retrieve, OCRonos provides a practical alternative for recovering usable text.
+Figure 4: Stacked histograms of probability density for qualitative evaluations of Common Corpus on a sample of 300,000 documents. Metric descriptions can be found in Appendix G.
+PII Removal.
+Personally Identifiable Information (PII), i.e., any information that can be used to distinguish or trace an individual’s identity, is protected under legislation such as GDPR.
+Consequently, the new regulations put restrictions on LLM training data.
+In large open datasets, there is a staggering amount of personal data in widely used datasets, e.g., large quantities of phone numbers in RedPajama, email addresses in S2ORC and peS2o, and IP addresses in the Stack (Elazar et al., 2024).
+To identify and replace PII, we use Microsoft’s Presidio (Mendels et al., 2018), an open-source state-of-the-art tool.
+With Presidio, we filtered out phone numbers, email addresses, IBANs, IP addresses, and URLs.
 With the base settings, Presidio identified on average 55-60% of texts that included phone numbers due to different possible number formats.
@@
 These transformations substantially alter the format of PII, which could undermine the model’s understanding of the text or interfere with its ability to process text with real PII.
-Instead, we replace PII with fictuous but realistic values.
-4.5 Toxicity Detection
+Instead, we replace PII with fictitious but realistic values.
+Deduplication.
+Our early experiments showed a negligible rate of duplication, which we attribute to the initial data curation: large institutions are incentivized to avoid re-digitizing the same texts.
+We also filtered out duplicates based on PDF metadata and used deduplicated sources wherever possible.
+Toxicity Detection.
 In addition to posing legal and regulatory issues, web data is a major source of harmful and biased content (Common Crawl was shown to contain sexual content, hate speech, and racial and gender biases (Luccioni and Viviano, 2021)) and often suffers from low-quality and machine-generated text (Dodge et al., 2021).
-Public Domain data, e.g., in Open Culture, do not pose the same legal challenges but introduce new ones.
-Many texts there are historical periodicals and monographs from at least 80 years ago.
-Cultural norms surrounding the discussion of certain ethnic groups, women, and themes such as violence have changed dramatically.
-Many of these texts, therefore, do not meet modern ethical standards.
-Training models on these texts would lead to the reproduction and circulation of harmful language.
-[…]
-5 Impact
-As early as August 2024, the first version of Common Corpus was highlighted as a “massive” part of pre-training data commons in European languages, especially contributing to making French a high-resource language (Ali and Pyysalo, 2024).
-In February 2025, Common Corpus was acknowledged as one of the main deliverables of the Paris AI Summit, as the “largest open database to train Large Language Models” [^15] https://www.elysee.fr/admin/upload/default/0001/17/373d6bcd8ea7f84a701c050bb9cab ccf3ed95c2b.pdf.
-Common Corpus has already been incorporated in a wide range of LLM projects.
-As of May 2025, seven European SLMs and LLMs have been trained on at least some part of Common Corpus, and more will come throughout the year 2025.
-Among these are the models in the Pleias 1.0 family, including three SLMs (350 million, 1.2 billion, and 3 billion parameters), which were exclusively trained on Common Corpus.
-As such, they are fully reproducible in Europe.
-They were trained on the Jean Zay supercomputer in France (using H100 GPUs, under a Grand Challenge compute grant), and, specifically for the 1.2B model on the TractoAI’s cloud platform [^16] https://tracto.ai/.
-They come with a new tokenizer with 65,536 unique tokens, also trained on a representative sample of the Common Corpus for increased training efficiency and better multilingual support.
-During training, we experimented with different data mixtures of Common Corpus and derived a “refined” version for late training of about 1 trillion tokens, with more intensive, curation-based filtering.
-Two models derived from Pleias 1.0, Pleias-RAG-350M and Pleias-RAG-1B, are currently leading evaluations in their size range for retrieval augmented evaluation.
-Part of the performance can be attributed to the solid coverage of the main European language in Common Corpus, as well as a very extensive proportion of PDF-extracted sources that are very close to content commonly processed in the RAG production setting.
-In January 2025, Barcelona Supercomputing Center released the Salamandra series of models under three different sizes: 2, 7, and 40 billion parameters (Gonzalez-Agirre et al., 2025).
-A major objective was enhanced support for a wide range of European languages, including lower-resource languages.
-The French part of Open Culture, French-PD, was incorporated into the overall data mixture.
-It was the main French dataset and the second-largest curated source.
-A French 7 billion parameters SLM, Lucie, incorporated a wider range of French sources from Common Corpus, not only the cultural heritage part but also some of the administrative sources (Gouvert et al., 2025).
-In addition, NeKo, a generalist language correction model trained by Nvidia, used the OCR-corrected part of Common Corpus both for training and evaluation (Lin et al., 2024).
-Anthropic also used Common Corpus in their experiments for feature visualization (Ameisen et al., 2025).
-Common Corpus is also being used to create new datasets.
-For instance, YouTube Commons was used to create two major multimodal datasets: FineVideo (43000 YouTube videos) [^17] https://huggingface.co/datasets/HuggingFaceFV/finevideo and Mosel (1 million audio hours, half from YouTube Commons) (Gaido et al., 2024), which was used to train the FAMA series of speech foundational models (Papi et al., 2025).
+Public Domain data, such as that in Open Culture, comprise historical periodicals and monographs from at least 80 years ago.
+As cultural norms have changed dramatically, many of these texts do not meet modern ethical standards.
+Training language models on these texts would lead to the reproduction and circulation of harmful language.
+To address this, we developed a pipeline to filter the public domain training data.
+We created a multilingual toxicity classifier, Celadon, a DeBERTa-v3-small model ($\sim$140M parameters), trained from scratch on 2M synthetically annotated samples.
+Celadon and the training dataset were released as parts of a separate work (Arnett et al., 2024).
+6 Evaluation
+We evaluate a sample from Common Corpus using various qualitative metrics and present the distributions in Figure 4 and provide details on metrics in Appendix G.
+Most data objects present high peaks at similar proportions in the expected range.
+Specific collections exhibit expected deviations: code data has a generally higher proportion of repetitions and a lower proportion of whitespace characters due to strict syntax and the presence of punctuation, while Open Government data presents lower linguistic diversity due to fixed terminology and expressions.
+We train two models on Common Corpus: PleIAs 350M and PleIAs 1.2B [^5] Repositories: [Uncaptioned image] PleIAs/Pleias-350m-Preview   [Uncaptioned image] PleIAs/Pleias-1.2b-Preview.
+The architecture is based on Llama.
+We train a custom Llama-style tokenizer with a vocabulary size of 65536 on a representative subset of Common Corpus.
+PleIAs 350M is trained on a filtered subset of Common Corpus, amounting to approximately 1T tokens.
+PleIAs 1.2B is trained on Common Corpus and three epochs of the filtered subset.
+The models were trained for 2944 and 23040 H100 hours, respectively.
+We evaluate our models on MultiBLiMP (Jumelet et al., 2025), XStoryCloze (Lin et al., 2022), and XCOPA (Ponti et al., 2020; see Table 2 for aggregated scores and Appendix G for per-language details).
+All evaluations were run using the LM Evaluation Harness (Biderman et al., 2024).
+Our models perform comparably to models trained on closed or non-permissively licensed data, and show outstanding performance on MultiBLIMP, which has more languages compared to other benchmarks.
+This is especially notable for our 350M model, which we compare to bigger models; it also outperforms models from the 1B range, except for Gemma 3 1B.
+Our models stably outperform OLMo 1B, which was also pre-trained on a publicly released dataset.
+Model Ours Gemma 3 XGLM BLOOM Ours Gemma 3 XGLM OLMo
+350M 270M 564M 560M 1.2B 1B 1.7B 1B
+MultiBLiMP 0.774 0.762 0.711 0.683 0.797 0.799 0.710 0.699
+XStoryCloze 0.509 0.533 0.537 0.532 0.526 0.594 0.569 0.517
+XCOPA 0.533 0.544 0.550 0.541 0.541 0.593 0.574 0.518
+Table 2: Benchmarking results. “Ours” refers to PleIAs models pre-trained on Common Corpus.
 7 Conclusion
-Common Corpus is a significant contribution to the open LLM research community.
-Since the release of the first version, the data from Common Corpus has been used in numerous LLM projects by several European labs and industry leaders like Anthropic.
 Through the release of Common Corpus and this paper with thorough documentation of data collection and curation, we show that LLM development is possible while strictly adhering to the regulatory norms.
+While Common Corpus is only large enough to train small models currently, the tools and methods we used to identify and curate the data may be used to expand the amount of permissively licensed open data.
 We hope that Common Corpus will grow as a critical infrastructure for open science LLM research and development and inspire future initiatives in the open.
 Limitations
-Common Corpus is far from collecting the whole range of available open data, which we call the open data paradox.
-Continuing the work, we are currently collecting more data in European languages as part of the LLMs4EU project [^18] https://www.openaire.eu/llms4eu.
-In addition to this, even though the language diversity of the data is large, the coverage for low-resource languages is still shallow.
-We are developing in this direction, planning for enlargement of the corpora in African languages (e.g., Swahili, Wolof, Bambara).
-Furthermore, the amount of collected data (2 trillion tokens) is suitable for pre-training of models of limited size, as in the Pleias model family (see Section 5).
-Larger models, however will require significantly larger amounts of data.
-Therefore, future collection of permissible data is highly encouraged by this work.
+Common Corpus is far from collecting the whole range of available open data, which we described as the open data paradox.
+Therefore, the future collection of permissive data is highly encouraged by this work.
+Furthermore, the collected amount of data (2 trillion tokens), when used alone, as our own small language model family (see Section 6), is suitable for pre-training of models of limited size, while larger ones require significantly larger amounts of data.
 In addition, Common Corpus naturally does not contain data for instruction-tuning and any forms of specialized tasks.
 Therefore, it is not directly suitable for task-specific fine-tuning.
-However, due to the multilingual, temporal, and semantic diversity of data, Common Corpus opens the opportunities for the creation of ethical fine-tuning datasets, as we did for the RAG models (see Section 5).
+However, due to the multilingual, temporal, and semantic diversity of data, Common Corpus opens the opportunities for the creation of ethical fine-tuning datasets.
 In Section 5, we described the tools we used for the data curation, filtering, and editing.
-Even though we used these methods responsibly and mitigated many issues overlooked by the counterparts (e.g., with toxicity detection and PII removal), none of the curation methods could naturally facilitate a hundred-percent accuracy.
+Although we used these methods responsibly and mitigated many issues overlooked by the counterparts (e.g., with toxicity detection), none of the curation methods could naturally facilitate a hundred percent accuracy.
 However, some issues, like OCR errors, present considerable challenges to the models and might even account for better handling of typos in the future.
 We would also like to mention that each data object is accompanied by sufficient metadata, and, if desired, LLM practitioners are free to filter out collections that might contain potential issues (as described in Section 5).
-Acknowledgments and Disclosure of Funding
+Acknowledgements
 Common Corpus is part of the Current AI initiative to accelerate the development of the building blocks of open AI — notably data — that serves and is shaped by the public interest.
@@
 The collection of the corpus has been largely facilitated thanks to major organizations committed to an open science approach for AI, namely AI Alliance, Mozilla, HuggingFace, Occiglot, and Eleuther AI.
-Appendix A Development
-Common Corpus was first conceptualized in May 2023 as a French corpus for the French public services.
-The initial aim was to create a dataset of 100-140B tokens and use that data to train a reproducible small language model, modeled on the Pythia suite (Biderman et al., 2023).
-The initial nucleus of Common Corpus was French-PD a large collection of 85B words from monographs and newspapers in the public domain [^19] Released as https://huggingface.co/datasets/PleIAs/French-PD-Books and https://huggingface.co/datasets/PleIAs/French-PD-Newspapers..
-A collection of various open datasets was first published under the name Common Corpus in March 2024202020Release blogpost: https://huggingface.co/blog/Pclanglais/common-corpus.
-At the time, it was a multilingual expansion of French-PD, retaining the initial focus on cultural heritage data but containing now about 500B words in the public domain from a variety of open collections in English, German, French, Italian, Spanish and other European languages.
-Throughout 2024, additional official releases gradually expanded this initial pretraining commons beyond the public domain, including Youtube Commons (in April 2024) and Finance Commons (in July 2024).
-In November 2024, Common Corpus became an unified corpus, through the convergence of newly collected datasets by PleIAs and pre-existing ones (especially Wikimedia Projects).
-It contained two trillon tokens with the current division in five collections [^21] Release blogpost: https://huggingface.co/blog/Pclanglais/two-trillion-tokens-open.
-The current version of Common Corpus, the one we describe in this paper, was released in February 2025 for the Paris AI Summit.
-Common Corpus will remain an evolving resource reflecting the need for a continuous open infrastructure rather than a fixed dataset, with HuggingFace ensuring a proper archiving of past versions for reproducibility.
-A primary driver of Common Corpus has been what we term the open data paradox: most of the content available under free licenses or uncopyrighted is not integrated into the existing pretraining datasets.
-Since GPT-3, the pretraining data infrastructure has been extensively reliant on web archiving and the constraints that stem from it.
-which include the lack of available tools for PDF hosting and processing, or for diverse language detection.
-The open data paradox had the effect of making the open pretraining ecosystem viable in the first place.
-In particular, the gradual filtering of web archives according to more and more ethical rules leaves only a tiny residual amount under open licenses (less than 2% of the original set).
-Appendix B Provenance
-B.1 Open Govenment
+Appendix A LLM Usage Statement
+In the process of developing this work, we utilized LLMs for grammar correction and occasionally as a rewriting tool. In addition, we involved LLMs in the process of data visualization.
+Appendix B Corpus Composition
+The documents in Common Corpus have the following structure:
+  • identifier: unique text identifier. In many cases, this also links to the original resource.
+  • open_type: one of the six leading collection groupings.
+  • collection: name of the sub-collection (Appendix D).
+  • license: details on the content licensing.
+  • date: date of creation of the resource, where known.
+  • title: title of the resource, when known, alternatively, the filename.
+  • creator: institution that published, collected, or curated the resource.
+  • language: automatically identified language.
+  • word_count: number of space-delimited words.
+  • token_count: number of tokens as calculated by PleIAs tokenizer.
+  • text: full text, without formatting.
+In Table 3, we present the token, word, and document counts for the Common Corpus collections.
+Dataset Documents Words Tokens
+Open Government 075,652,998 0,257,561,830,682 0,407,067,554,189
+Open Culture 093,156,602 0,549,608,763,966 0,885,982,490,090
+Open Science 019,220,942 0,147,305,783,453 0,281,193,563,789
+Open Code 202,765,051 0,077,669,169,092 0,283,227,402,898
+Open Web 096,165,348 0,033,208,509,065 0,073,217,485,489
+Open Semantic 030,072,707 0,023,284,201,782 0,067,958,671,827
+Total 517,033,648 1,088,638,258,040 1,998,647,168,282
+Table 3: Dataset composition of Common Corpus. For each collection, we report the total number of documents, words (whitespace-separated), and tokens.
+In Table 4, we present the top ten licences of Common Corpus documents.
+License type Tokens
+Public Domain 1,138,508,375,958
+$C$-By 0,287,749,264,457
+MIT 0,142,694,227,607
+$C$-By-SA 0,074,768,060,836
+Apache-2.0 0,068,750,977,037
+BSD-3-Clause 0,018,483,944,333
+Open license 0,010,432,513,767
+BSD-2-Clause 0,005,497,145,480
+$C$-BY-4.0 0,002,110,966,243
+CC0-1.0 0,001,877,206,195
+Table 4: Token counts for the ten most common licenses in Common Corpus.
+In Table 5, we present the top-50 languages in Common Corpus by token count.
+The token counts are presented in terms of our BPE tokenizer used to train the models described in Section 6, which was trained on a representative subsample of Common Corpus.
+To verify that our tokenizer serves as a strong baseline for token counts, we show its fertility in Appendix C.
+Language Documents Words Tokens
+English 154,175,907 634,794,970,595 968,757,721,747
+French 035,245,624 162,061,620,874 275,358,437,630
+German 011,385,377 056,674,819,173 112,127,458,251
+Spanish 006,530,094 026,215,767,271 046,514,142,421
+Latin 002,367,110 016,189,444,325 036,031,591,540
+Italian 003,804,052 013,207,129,356 024,681,637,575
+Polish 002,640,613 005,086,555,167 012,146,688,669
+Greek 000844,122 003,796,018,483 011,376,498,056
+Portuguese 001,756,922 005,234,373,473 010,262,747,943
+Russian 002,762,818 003,222,919,854 009,439,453,633
+Dutch 003,206,382 003,791,928,728 008,058,934,080
+Danish 002,270,459 002,840,121,206 006,941,827,931
+Slovak 000683,174 002,320,831,403 005,148,967,838
+Czech 000946,534 001,966,829,784 004,798,558,092
+Indonesian 001,023,361 001,660,129,567 004,381,878,823
+Estonian 000685,180 001,613,093,565 004,379,534,617
+Hungarian 000888,780 001,527,981,866 004,110,878,972
+Swedish 003,250,289 001,782,642,556 004,014,927,806
+Finnish 000931,201 001,356,653,003 003,943,036,413
+Maltese 000480,491 001,421,372,608 003,646,102,921
+Bulgarian 000576,997 001,444,748,621 003,422,182,324
+Lithuanian 000539,906 001,192,586,834 003,097,400,907
+Romanian 000725,766 001,398,178,029 002,909,452,579
+Japanese 001,409,956 000204,698,439 002,738,872,745
+Arabic 001,291,439 000827,392,227 002,682,255,180
+Slovenian 000504,492 001,107,467,643 002,602,943,380
+Latvian 000364,503 000966,005,785 002,579,350,119
+Ukrainian 001,378,390 000786,829,868 002,561,253,212
+Croatian 000443,669 000984,577,685 002,400,762,140
+Chinese 001,426,017 000329,649,628 002,238,230,225
+Haitian Creole 000292,718 000762,225,686 001,420,886,048
+Turkish 000572,633 000380,551,359 001,206,732,266
+Cebuano 006,123,694 000598,016,557 001,080,404,897
+Norwegian Nynorsk 001,026,745 000405,911,677 001,036,914,970
+Irish 000139,815 000371,175,864 000958,688,647
+Castilian 000050,785 000457,490,798 000893,080,621
+Serbian 000698,037 000292,399,267 000822,519,328
+Hebrew 000343,406 000227,869,199 000763,029,488
+Catalan 000803,532 000368,144,006 000736,220,789
+Korean 000653,460 000143,729,419 000677,716,397
+Persian 000983,632 000209,078,446 000668,287,975
+Vietnamese 001,296,789 000267,190,147 000593,203,801
+Norwegian 000628,798 000252,133,513 000561,694,623
+Armenian 000306,959 000109,852,686 000539,900,475
+Hindi 000210,984 000127,830,222 000463,832,707
+Yiddish 000059,158 000122,100,684 000463,526,449
+Welsh 000309,573 000191,827,552 000438,133,529
+Occitan 000288,195 000195,380,666 000357,267,635
+Georgian 000172,532 000071,678,084 000349,884,144
+Basque 000439,582 000139,077,600 000348,891,265
+Table 5: Top-50 languages in Common Corpus by token count. Each language is presented with its number of documents, words, and tokens in the corpus. The rows are ordered by the token count.
+Appendix C Tokenizer Details
+Figure 5: Comparing the fertility of PleIAs tokenizer (marked as “Ours 1B”) and other language models for six languages.
+The data source for all languages is the devtest set of Flores200 (Costa-jussà et al., 2022).
+In Figure 5, we show how our tokenizer with a vocabulary of size 65,536, trained on a representative subsample of Common Corpus, compares to other language model tokenizers.
+Our tokenizer is outperformed only by Gemma 3, which has a tokenizer four times larger.
+Appendix D Provenance
+D.1 Open Govenment
 In this section, we describe the provenance and present token counts and main languages for the two sub-collections of Open Government: Finance Commons and Legal Commons.
-B.1.1 Finance Commons
+D.1.1 Finance Commons
 Dataset Main Languages Documents Tokens
@@
 Table 6: Finance Commons sources distribution with languages.
-The datasets that make up Finance Commons [^22] https://huggingface.co/collections/PleIAs/finance-commons-66925e1095c7fa6e6828e26c are presented in Table 4.
-Here, we also present the provenance details for each of the parts of Finance Commons:
+The datasets that make up Finance Commons are presented in Table 6. Here, we also present the provenance details for each of the parts of Finance Commons:
   • Securities and Exchange Commission (SEC).
@@
   • World Trade Organization (WTO).
-This dataset comprises documents from WTO’s official Documents Online platform [^25] https://docs.wto.org/dol2fe/Pages/FE_Search/FE_S_S005.aspx.
+This dataset comprises documents from WTO’s official Documents Online platform.
 The documents cover the years 1995 to 2024.
@@
 Some documents are available in other languages, e.g., Chinese, Korean, Arabic, German, and Portuguese.
-Also released separately as WTO-PDF [^26] https://huggingface.co/datasets/PleIAs/WTO-PDF.
+Also released separately as WTO-PDF.
   • French Authority for Financial Market (AMF).
@@
 The documents are primarily in French.
-Also released separately as AMF-PDF [^28] https://huggingface.co/datasets/PleIAs/AMF-PDF.
+Also released separately as AMF-PDF.
   • Tenders Electronic Daily (TED) EU Tenders.
@@
 This dataset comprises documents from GATT, which was an organization that promoted international commerce and the reduction of trade barriers among member states.
-Public documents were made available by the General Council of the WTO in 2006303030https://www.wto.org/english/docs_e/gattdocs_e.htm.
+Public documents were made available by the General Council of the WTO in 2006101010https://www.wto.org/english/docs_e/gattdocs_e.htm.
 The documents span from January 1, 1946, to September 6, 1996.
 Most of the documents are in English, but there are also documents in French, Spanish, and other languages.
-B.1.2 Legal Commons
+D.1.2 Legal Commons
 Dataset Languages Tokens
@@
 Here, we present the provenance details for each of the parts of Legal Commons:
-B.2 Open Culture
+D.2 Open Culture
 Corpus Language Domain Tokens
@@
 Arabic PD Arabic Books 000.3B
-Table 6: Subsets of Open Culture with language coverage, domains, and token counts.
+Table 8: Subsets of Open Culture with language coverage, type of document, and token count.
 Large portion of data in Open Culture part of the Common Corpus was built on top of the following collection-as-data initiatives:
@@
 The texts were originally published between 1737 and 1905.
-The corpus was originally digitized between 2009 and 2011.
-The corpus was made available by the Institut für Deutsche Sprache in 2013.
+The corpus was originally digitized between 2009 and 2011 and made available by the Institut für Deutsche Sprache in 2013.
   • Spanish PD Books.
@@
 To ensure that these texts are in the public domain, we have retained exclusively titles published prior to 1884.
-B.3 Open Science
-In Table 7, we present the total token counts per collection inside of the Open Science part of Common Corpus.
+D.3 Open Science
+In Table 9, we present the token counts inside of the Open Science collections.
 Dataset Tokens
@@
 Table 9: Token count by dataset Open Science.
-B.4 Open Code
+D.4 Open Code
+Table 10 shows the number of tokens for the top ten coding languages and frameworks in Open Code.
 Language Tokens
@@
 Table 10: Token counts by programming language or framework.
-Table 8 shows the number of tokens for the top ten coding languages and frameworks in Open Code.
-Appendix C Open Culture Verification
+Appendix E Open Culture Verification
 Here, we describe the rights verification process that we applied for cultural data objects:
@@
 14) and common practice among GLAM reusers like Wikimedia Commons, we consider that the simple act of digitization does not provide any additional rights.
-Appendix D Cleaning and Curation
-D.1 Text Segmentation
+Appendix F Cleaning and Curation
+We release the cleaning and curation tools as part of our Bad Data Toolbox, a suite designed to make challenging digitized resources usable for downstream NLP tasks and large-scale corpus construction.
+F.1 Text Segmentation: Segmentext
+Repository: [Uncaptioned image] PleIAs/Segmentext
+Segmentext is a token classification model trained for the structural segmentation of documents, with a specific focus on robustness to digitisation artefacts and broken or unstructured input text.
+In contrast to layout-based segmentation approaches that operate on visual signals such as bounding boxes or font information, Segmentext works exclusively on raw character sequences, reconstructing editorial structure from text alone.
+This makes it applicable in post-OCR pipelines where layout information has been lost.
+The model is a DeBERTa-v2 (He et al., 2021) variant for token classification, fine-tuned on 3,500 manually annotated documents.
+The training data was drawn from three datasets spanning complementary domains: Open Government, Open Culture, and Open Science.
+This domain diversity is intended to give the model broad applicability across document types and languages.
+Segmentext supports the following text segmentation:
+  • Text: the text itself.
+  • Title: any type of heading.
+  • Table: tabular content.
+  • Separator: a segmentation separator, mostly based on newline, with some variations due to text segmentation understanding.
+  • Dialog: any kind of speaker-attributed intervention.
+  • Bibliography: statement of a specific bibliographic reference, either in a bibliography section or a footnote.
+  • Contact: personal information, which can be especially useful in the context of PII removal.
+  • Paratext: any non-meaningful text included in standard documents like headers, page numbering, section recall, etc.
+  • Author: author names and signatures.
+  • Date: statement of date and time, common in letters and newspaper articles.
+  • Keyword: list of keywords, especially common in scientific publications.
 Here is an example input text for the Segmentext model:
@@
 Following the previous year, the group continued to diversify its holdings into investment-grade corporate bonds.
-It should be noted that bonds and term loans are held to maturity in accordance with the group’s business model policy of "inflows".
+It should be noted that bonds and term loans are held to maturity in accordance with the group’s business model policy of ”inflows”.
 Technical liabilities on insurance contracts.
@@
 Income statement.
-The income and expenses recognized for the insurance contracts issued by the group appear in the income statement in "Net income of other activities" and "Net expense of other activities".
+The income and expenses recognized for the insurance contracts issued by the group appear in the income statement in ”Net income of other activities” and ”Net expense of other activities”.
 Risk management.
-The group adopts a "prudent approach" to its management of the risks to which it could be exposed through its insurance activities.
+The group adopts a ”prudent approach” to its management of the risks to which it could be exposed through its insurance activities.
 Risk of counterparty.
@@
 Example output:
-D.2 OCR Error Detection
-OCRoscope.
-To illustrate this approach, this long text is correctly identified as French with >99% confidence by cld2, as despite the many mistakes, there are enough non-ambiguous French words:
-NOUVELLES POLI TI QÛ E S.
+F.2 OCR Error Detection: OCRoscope
+Repository: [Uncaptioned image] PleIAs/OCRoscope
+OCRoscope is a lightweight Python library for estimating the quality of OCR-processed text without access to a ground-truth reference.
+Rather than relying on character- or word-level edit distance against a gold standard, OCRoscope exploits the differential behaviour of language detection models across text granularities.
+On long texts, language detection is robust to noise; on short character sequences, it degrades rapidly in the presence of OCR errors.
+The library leverages this property by applying language detection (via pycld2) to rolling 7-grams extracted from the input text and comparing the per-gram language assignments against the document-level language detected on the full text.
+The proportion of n-grams assigned to a mismatching or unknown language yields the OCR quality score: a value close to 1 indicates high-quality text, while lower values signal increasing levels of digitisation noise.
+OCRoscope returns two primary metrics: a standardised OCR quality rate and a standardised rate of non-character content (punctuation, digits, whitespace anomalies).
+The tool supports over 80 languages via cld2’s language coverage.
+The choice of cld2 over more accurate successors such as cld3 is intentional: cld2’s greater sensitivity to short noisy sequences makes it a more discriminating signal for OCR quality estimation, yielding a score distribution that is well spread across the quality spectrum and therefore easier to threshold for downstream filtering decisions.
+The tool is designed for pipeline use cases in which large corpora must be rapidly triaged: it provides a fast, reference-free quality signal that can guide decisions about whether texts warrant OCR correction, manual review, or outright exclusion.
+Below we illustrate OCRoscope performance on this long text, which is correctly identified as French with >>99% confidence by cld2.
+Despite the many mistakes, there are enough non-ambiguous French words:
+NOUVELLES POLI TI QˆU E S.
 Suede.
@@
 11 paraît que k Uc-régeik a des craintes ; il a fait venir chez lji les membres c Ij“‘ tribunal 4e la cour , et leur a rtmis son lesfca n at.
-La fermentation qu’a causée 1 ,’ari r?tavh n k M p v riote Thorild tî’est pas appaisée y le luigage qv’il a yailé an duc-régent a été bien entendu par le peu) k y ir M» (U i n’entendrait pas l’apostrdphe suivante ?
-ttRxc3xa7nd >la libuk à r otre raison , et ne et nous force pas de i’ache’ef r i te n :e sang,.
+La fermentation qu’a causée 1 ,’ari r?tavh n k M p v riote Thorild tî’est pas appaisée y le luigage qv’il a yailé an duc-régent a été bien entendu par le peu) k y ir M (U i n’entendrait pas l’apostrdphe suivante ?
+ttRxc3xa7nd >>la libuk à r otre raison , et ne et nous force pas de i’ache’ef r i te n :e sang,.
 Le duc a fait x,épa4idre sur-le-champ une fjtbprijuun à te us les habitans di\$ Toyaume , pour les detourntr de mr laisser sé luire par de fa,ux bruits et des jugemens pe rver\$ , e i en même temps l’ordre a.
@@
 (Mercure Français, 1793, January 25th)
-Yet one short n-gram ("n k M p v riote Thorild") is classified as unknown by cld2.
-OCRerrcr.
-The following is a low-error example sentence taken from Common Corpus:
+Yet one short n-gram (“n k M p v riote Thorild”) is classified as unknown by cld2.
+Complete processing by OCRoscope yielded a 41% rate of non-recognized 7-grams, resulting in an OCR quality rate of 59%.
+In comparison, the self-estimated rate of OCR-validated words by the French National Library is significantly higher (85%) for the whole document.
+F.3 OCR Error Detection: OCRerrcr
+Repository: [Uncaptioned image] PleIAs/OCRerrcr
+OCRerrcr is a small token classification model trained to identify individual OCR errors at the token level within a text.
+Unlike OCRoscope, which produces a document-level quality score, OCRerrcr generates a fine-grained annotation of which tokens are likely to be erroneous, making it suitable for targeted correction workflows and for producing spatially explicit error rate estimates.
+The model is a DeBERTA-v3 (He et al., 2023) with 400M parameters fine-tuned on a corpus of approximately 1,000 documents with manually labelled OCR errors, drawn from Open Government and Open Culture.
+This dual-domain training is intended to give the model coverage across both modern administrative language and historical or literary language, which present different error profiles.
+OCRerrcr provides an accurate reference-free OCR error rate estimate, outperforming OCRoscope in precision, particularly on documents with a low or moderate error rate.
+However, it scales less efficiently than OCRoscope to very large corpora due to its per-token inference cost.
+The model’s name is itself a playful illustration of the problem it addresses: OCRerrcr is a plausible OCR misreading of OCRerror, with the vowels dropped as they might be in a degraded scan.
+The following is a low-error example sentence taken from Open Culture:
 They did not approach cer, but turned away and passed irom her presence, filled with sorrow and moved with sympathy, which her intense emotions seemed to communicate to even these thoughtless young men of the tho plains.
-And the OCRerrcr detection (with formatting for clarity):
-They did not approach <<<er>>>cer,<<</er>>> but turned away and passed <<<er>>>irom<<</er>>> her presence, filled with sorrow and moved with sympathy, which her intense emotions seemed to communicate to even these thoughtless young men of the <<<er>>>tho<<</er>>> plains.
-D.3 OCR Correction
-Here is an example of text containing various OCR errors:
+The version with the errors highlighted by OCRerrcr is presented below:
+They did not approach <<er>>cer,<</er>> but turned away and passed <<er>>irom<</er>> her presence, filled with sorrow and moved with sympathy, which her intense emotions seemed to communicate to even these thoughtless young men of the <<er>>tho<</er>> plains.
+F.4 OCR Error Correction: OCRonos
+Repository: [Uncaptioned image] PleIAs/OCRonos
+OCRonos is a generative language model trained for the correction of badly digitised text, including OCR errors, incorrect word splits and merges, and broader structural artefacts introduced during digitisation.
+It belongs to the corrective tier of the Bad Data Toolbox, operating downstream of quality detection tools such as OCRoscope and OCRerrcr.
+The training data for OCRonos consists of a diverse multilingual set of OCR-processed texts drawn from Open Culture and Open Government.
+The model is fine-tuned from Llama-3-8B (Grattafiori et al., 2024).
+OCRonos is designed to be conservative: it is generally faithful to the original source material and rarely rewrites words that are not erroneous.
+On highly degraded inputs, however, it can function as a synthetic rewriting tool rather than a strict corrector.
+A known issue at scale is the occasional inclusion of spuriously repeated words, which can be filtered in post-processing.
+The model also mitigates a common failure mode observed in smaller generalist LLMs, namely language switching: faced with heavily noisy input, models such as GPT-3.5 or Claude Haiku have been observed to drift into a different language or script during correction.
+OCRonos is specifically trained to resist this behaviour.
+Below we present an exemplar text containing various OCR errors:
 Theguaran tees offered cover death,disability,r e dundancy andunem ployment aspartof aloanprotect ion insurance policy.
 These types o f risk are controlled throu ghthe use o f app ropriate morta litytables,statistica lchecksonloss rat ios for thepopulation groups insure dandthrough ar e insurance program.
-And here is the text corrected by our model, OCRonos:
+The following is the version corrected by OCRonos:
 The guarantees offered cover death, disability, redundancy, and unemployment as part of a loan protection insurance policy.
 These types of risk are controlled through the use of appropriate mortality tables, statistical checks on loss ratios for the population groups insured, and through the insurance program.
+F.5 Toxicity Detection: Celadon
+Repository: [Uncaptioned image] PleIAs/celadon
+Celadon is a multilingual toxicity classifier designed specifically for the challenges posed by public-domain pre-training data: historical language, OCR noise, and multilingual content spanning nine languages.
+It was developed as part of the Toxic Commons project, described in Arnett et al.
+(2024), and was applied to filter harmful content in Common Corpus prior to its release.
+The model is a fine-tune of DeBERTa-v3-small (He et al., 2023) with five independent classification heads, each corresponding to one dimension of toxicity: race and origin-based bias (including racism, xenophobia, and bias against immigration status), gender and sexuality-based bias (including sexism, homophobia, and transphobia), religious bias, ability bias (bias relating to physical, mental, or intellectual disability), and violence and abuse (graphic violence, threats, and incitement).
+Each head is a separate linear layer with a custom weighted cross-entropy loss function to handle class imbalance.
+The model was trained on 600,000 samples drawn from ToxicCommons, a dataset of approximately 2 million annotated public-domain text segments from 13 multilingual cultural heritage collections.
+Annotations were generated in two stages: a small set of human annotations was used to calibrate a prompt for the instruct version of Llama 3.1-8B (Grattafiori et al., 2024), which then produced annotations at scale across five toxicity dimensions using a 0–3 severity scale.
+The training set was balanced to equalize non-toxic and toxic samples.
+Celadon’s key design rationale is efficiency: it achieves toxicity detection performance comparable to Llama 3.1-8B-Instruct while being over 40 times faster (5 minutes versus 3.4 hours per 100,000 samples), making it practical for large-scale corpus filtering.
+The model is optimised for the specific challenges of its target domain: historical text, OCR noise, and non-English content, and may not generalise without adaptation to web-text corpora, for which it was not designed.
+Appendix G Evaluations
+Figure 4 presents the distribution of six document-level quality and character composition metrics across the five open types in a 300,000-document sample of Common Corpus:
+  • Character Repetition measures the proportion of character pairs belonging to repeated character bigram sequences within a document.
+This value is expected to be higher for code data, due to specific syntax and varied punctuation.
+  • Type-Token Ratio is the ratio of unique word forms to the total number of tokens.
+It captures lexical diversity; documents with values close to 1 are typically very short, while low values can indicate highly repetitive text.
+For Open Government, this value is expected to be lower as regulatory texts usually contain fixed, distinct terms and repetitive formulations.
+  • Top Word Proportion is the share of all tokens accounted for by the single most frequent word in a document. This is another measure of linguistic difersity.
+  • Alphanumeric Characters denotes the fraction of characters that are letters or digits. This proportion might be lower in code due to reliance on punctuation and whitespace symbols.
+  • Whitespace Characters gives the proportion of whitespace characters, which tends to be higher in source code and tabular data than in running prose.
+  • Uppercase Characters measures the share of uppercase letters and separates Open Source content, which is densely populated with identifiers, constants, and type annotations, from Open Culture and Open Government documents that follow standard typographic conventions.
+In Tables 111213, and 14, we present per-language scores for the studied benchmarks.
+On MultiBLIMP, most of the scores are significantly above random (0.5); therefore, we also highlight the best and second-best scores.
+Model Ours Gemma 3 XGLM BLOOM Ours Gemma 3 XGLM OLMo
+350M 270M 564M 560M 1.2B 1B 1.7B 1B
+abk 0.550 0.750 0.475 0.525 0.675 0.675 0.325 0.800
+aln 0.733 0.755 0.709 0.700 0.728 0.750 0.675 0.690
+amh 0.946 0.929 0.911 0.973 1.000 0.955 0.848 0.964
+apu 0.964 0.964 0.893 0.786 0.964 0.929 0.964 0.893
+aqz 0.214 0.357 0.429 0.500 0.429 0.429 0.714 0.214
+arb 0.877 0.913 0.895 0.923 0.900 0.951 0.887 0.782
+azz 0.734 0.744 0.729 0.720 0.729 0.758 0.773 0.686
+bel 0.799 0.795 0.574 0.608 0.853 0.896 0.577 0.611
+ben 0.571 0.762 1.000 0.810 0.762 0.905 0.857 0.524
+bho 0.676 0.647 0.588 0.588 0.706 0.794 0.618 0.588
+bor 0.722 0.631 0.697 0.610 0.697 0.627 0.680 0.668
+bre 0.942 0.815 0.554 0.604 0.938 0.946 0.615 0.685
+bua 0.680 0.718 0.670 0.718 0.670 0.641 0.699 0.660
+bul 0.872 0.880 0.969 0.623 0.897 0.945 0.976 0.735
+cat 0.885 0.852 0.961 0.950 0.919 0.931 0.953 0.735
+ces 0.824 0.808 0.579 0.597 0.858 0.891 0.603 0.668
+chu 0.670 0.648 0.582 0.635 0.659 0.663 0.593 0.632
+cym 0.771 0.730 0.633 0.611 0.828 0.796 0.610 0.796
+dan 0.980 1.000 0.840 0.800 0.980 1.000 0.740 0.940
+deu 0.967 0.949 0.961 0.754 0.977 0.981 0.969 0.886
+egy 0.409 0.409 0.409 0.455 0.500 0.455 0.409 0.455
+ell 0.931 0.937 0.985 0.676 0.948 0.975 0.984 0.842
+eng 0.981 0.979 0.973 0.960 0.983 0.987 0.974 0.984
+est 0.729 0.699 0.885 0.561 0.800 0.800 0.915 0.587
+eus 0.916 0.927 0.963 0.952 0.916 0.938 0.982 0.905
+fao 0.707 0.647 0.509 0.556 0.772 0.806 0.552 0.681
+fas 0.756 0.810 0.567 0.577 0.837 0.919 0.565 0.655
+fin 0.736 0.744 0.947 0.562 0.809 0.893 0.935 0.645
+fra 0.994 0.963 0.963 0.984 0.993 0.989 0.976 0.928
+frm 0.997 0.741 0.745 0.905 0.997 0.847 0.820 0.765
+fro 0.782 0.701 0.686 0.709 0.822 0.725 0.694 0.679
+gla 0.955 0.924 0.924 0.939 0.909 0.924 0.939 0.970
+gle 0.750 0.821 0.750 0.786 0.679 0.750 0.750 0.714
+glg 0.849 0.807 0.798 0.788 0.879 0.895 0.789 0.754
+got 0.630 0.642 0.588 0.599 0.645 0.631 0.580 0.598
+grc 0.824 0.719 0.683 0.623 0.887 0.758 0.711 0.707
+guj 1.000 1.000 1.000 1.000 1.000 0.857 1.000 1.000
+hbo 0.711 0.712 0.683 0.706 0.737 0.764 0.629 0.679
+hbs 0.892 0.848 0.619 0.603 0.922 0.929 0.616 0.723
+heb 0.853 0.829 0.609 0.642 0.876 0.868 0.585 0.667
+hin 0.854 0.934 0.975 0.971 0.916 0.966 0.977 0.748
+hit 0.620 0.640 0.520 0.600 0.560 0.480 0.540 0.540
+hsb 0.683 0.677 0.624 0.629 0.667 0.694 0.624 0.591
+hun 0.928 0.867 0.728 0.692 0.938 0.925 0.686 0.740
+hye 0.883 0.898 0.631 0.623 0.929 0.946 0.662 0.671
+hyw 0.813 0.781 0.583 0.608 0.894 0.859 0.551 0.629
+isl 0.710 0.751 0.653 0.660 0.767 0.863 0.636 0.667
+ita 0.925 0.910 0.915 0.670 0.952 0.965 0.915 0.791
+Table 11: Multilingual benchmarking results on MultiBLIMP (ISO 639 language codes a*–i* in alphabetical order).
+“Ours” refers to PleIAs models pre-trained on Common Corpus.
+Within each model group, the best score is in bold, and the second-best is underlined.
+Model Ours Gemma 3 XGLM BLOOM Ours Gemma 3 XGLM OLMo
+350M 270M 564M 560M 1.2B 1B 1.7B 1B
+kat 0.931 0.951 0.917 0.760 0.951 0.951 0.809 0.907
+kaz 0.705 0.792 0.647 0.682 0.780 0.844 0.682 0.688
+kir 0.930 0.843 0.914 0.919 0.935 0.924 0.843 0.941
+kmr 0.710 0.662 0.588 0.579 0.761 0.748 0.577 0.588
+koi 0.628 0.488 0.605 0.558 0.558 0.651 0.628 0.605
+kpv 0.641 0.591 0.553 0.547 0.700 0.628 0.581 0.547
+krl 0.650 0.612 0.688 0.573 0.612 0.642 0.704 0.573
+kxh 0.483 0.433 0.475 0.333 0.458 0.450 0.442 0.483
+lat 0.874 0.651 0.578 0.575 0.925 0.730 0.568 0.625
+lav 0.791 0.747 0.616 0.604 0.844 0.862 0.611 0.623
+lij 0.783 0.744 0.669 0.638 0.780 0.807 0.701 0.665
+lit 0.928 0.848 0.745 0.740 0.947 0.932 0.736 0.779
+mar 0.737 0.717 0.735 0.667 0.713 0.776 0.726 0.776
+mdf 0.537 0.622 0.524 0.537 0.622 0.585 0.561 0.500
+mkd 0.923 0.974 0.769 0.769 0.821 1.000 0.590 0.718
+myv 0.608 0.614 0.565 0.560 0.636 0.619 0.532 0.547
+nds 0.736 0.729 0.674 0.663 0.749 0.732 0.674 0.700
+nhi 0.526 0.579 0.474 0.632 0.553 0.579 0.447 0.500
+nld 0.924 0.912 0.620 0.627 0.954 0.963 0.663 0.829
+olo 0.679 0.668 0.795 0.611 0.753 0.711 0.842 0.595
+orv 0.733 0.721 0.690 0.636 0.757 0.744 0.707 0.667
+ota 0.879 0.929 0.899 0.848 0.939 0.949 0.889 0.828
+pcm 1.000 1.000 1.000 0.923 1.000 0.962 1.000 0.885
+pol 0.892 0.849 0.624 0.634 0.930 0.931 0.628 0.725
+por 0.948 0.933 0.939 0.955 0.965 0.972 0.920 0.872
+quc 0.779 0.672 0.649 0.740 0.740 0.664 0.679 0.656
+ron 0.868 0.874 0.638 0.608 0.903 0.928 0.640 0.793
+rus 0.921 0.916 0.937 0.727 0.952 0.963 0.954 0.819
+sah 0.688 0.771 0.736 0.792 0.708 0.681 0.701 0.764
+san 0.657 0.666 0.612 0.609 0.670 0.678 0.618 0.620
+slk 0.797 0.739 0.528 0.570 0.824 0.861 0.533 0.588
+slv 0.882 0.796 0.618 0.636 0.903 0.854 0.622 0.711
+sme 0.689 0.705 0.653 0.660 0.681 0.700 0.668 0.659
+sms 0.833 0.802 0.779 0.757 0.821 0.779 0.764 0.768
+spa 0.959 0.945 0.950 0.966 0.970 0.973 0.956 0.896
+sqi 0.786 0.823 0.494 0.588 0.823 0.881 0.539 0.765
+swe 0.995 0.995 0.970 0.950 0.995 1.000 0.985 0.990
+tam 0.942 0.942 0.969 0.966 0.932 0.953 0.976 0.746
+tpn 0.111 0.000 0.111 0.111 0.222 0.111 0.000 0.000
+ttc 0.478 0.478 0.493 0.449 0.449 0.464 0.435 0.478
+tur 0.766 0.804 0.829 0.700 0.814 0.908 0.857 0.716
+uig 0.764 0.760 0.722 0.732 0.755 0.757 0.686 0.740
+ukr 0.874 0.892 0.640 0.606 0.911 0.946 0.648 0.704
+urb 0.538 0.462 0.462 0.231 0.538 0.462 0.462 0.462
+urd 0.858 0.925 0.956 0.935 0.896 0.964 0.960 0.736
+uzb 0.900 0.880 0.780 0.720 0.900 0.940 0.760 0.880
+vep 0.572 0.588 0.588 0.481 0.631 0.626 0.567 0.524
+wbp 0.250 0.000 0.167 0.250 0.250 0.083 0.250 0.250
+wol 0.892 0.881 0.851 0.891 0.868 0.879 0.854 0.823
+xcl 0.679 0.674 0.585 0.636 0.711 0.702 0.613 0.616
+xnr 0.791 0.756 0.779 0.616 0.744 0.814 0.721 0.733
+xpg 0.820 0.900 0.820 0.860 0.880 0.900 0.900 0.900
+yrl 0.689 0.722 0.594 0.604 0.683 0.669 0.626 0.601
+Table 12: Multilingual benchmarking results on MultiBLIMP (ISO 639 language codes k*–y* in alphabetical order).
+“Ours” refers to PleIAs models pre-trained on Common Corpus.
+Within each model group, the best score is in bold, and the second-best is underlined.
+Model Ours Gemma 3 XGLM BLOOM Ours Gemma 3 XGLM OLMo
+350M 270M 564M 560M 1.2B 1B 1.7B 1B
+ar 0.475 0.492 0.500 0.521 0.477 0.572 0.525 0.473
+ca 0.514 0.513 0.567 0.561 0.535 0.600 0.602 0.509
+en 0.569 0.614 0.606 0.612 0.617 0.698 0.645 0.704
+es 0.520 0.558 0.549 0.555 0.543 0.628 0.593 0.556
+eu 0.516 0.524 0.531 0.538 0.514 0.531 0.561 0.503
+gl 0.490 0.486 0.461 0.467 0.532 0.576 0.484 0.459
+hi 0.509 0.541 0.520 0.549 0.515 0.598 0.557 0.493
+id 0.500 0.544 0.542 0.555 0.518 0.631 0.583 0.498
+my 0.492 0.507 0.515 0.475 0.498 0.518 0.537 0.475
+ru 0.503 0.547 0.562 0.488 0.531 0.634 0.600 0.512
+sw 0.500 0.507 0.531 0.500 0.510 0.551 0.563 0.494
+te 0.542 0.562 0.559 0.557 0.553 0.592 0.581 0.532
+zh 0.491 0.536 0.533 0.545 0.494 0.594 0.561 0.511
+Table 13: Multilingual benchmarking results on XStoryCloze. “Ours” refers to PleIAs models pre-trained on Common Corpus. Languages are represented as two-letter codes in ISO 639.
+Model Ours Gemma 3 XGLM BLOOM Ours Gemma 3 XGLM OLMo
+350M 270M 564M 560M 1.2B 1B 1.7B 1B
+es 0.566 0.590 0.604 0.620 0.614 0.678 0.664 0.524
+et 0.518 0.498 0.554 0.488 0.500 0.536 0.568 0.480
+eu 0.504 0.514 0.512 0.502 0.518 0.502 0.534 0.516
+ht 0.522 0.504 0.548 0.500 0.524 0.518 0.556 0.534
+id 0.534 0.578 0.574 0.596 0.558 0.690 0.646 0.544
+it 0.542 0.524 0.536 0.502 0.562 0.648 0.536 0.488
+qu 0.522 0.492 0.492 0.500 0.500 0.502 0.522 0.506
+sw 0.528 0.546 0.530 0.516 0.538 0.544 0.562 0.510
+ta 0.536 0.560 0.562 0.558 0.556 0.566 0.550 0.550
+th 0.546 0.538 0.550 0.538 0.550 0.584 0.580 0.532
+tr 0.530 0.566 0.544 0.528 0.542 0.606 0.536 0.530
+vi 0.550 0.598 0.584 0.602 0.526 0.694 0.630 0.494
+zh 0.536 0.564 0.554 0.588 0.544 0.640 0.584 0.522
+Table 14: Multilingual benchmarking results on XCopa. “Ours” refers to PleIAs models pre-trained on Common Corpus. Languages are represented as two-letter codes in ISO 639.